### v1

# On the Capacity of Self-Attention

Abstract
While self-attention is known to learn relations among tokens, we lack a formal understanding of its capacity: how many distinct relations can a single layer reliably recover for a given budget?
To formalize this, we introduce Relational Graph Recognition (RGR), where the key-query channel represents a graph on $m$ items with $m^{\prime}$ directed edges, and, given a context of items, must recover the neighbors of each item. We measure resources by the total key dimension $D_{K}=h\,d_{k}$. Within this framework, we analytically derive a capacity scaling law and validate it empirically. We show that $D_{K}=\Theta(m^{\prime}\log m^{\prime}/d_{\text{model}})$ is both necessary (information-theoretic lower bound) and sufficient (explicit construction) in a broad class of graphs to recover $m^{\prime}$ relations. This scaling law directly leads to a new, capacity-based rationale for multi-head attention that applies even when each item only attends to a single target. When embeddings are uncompressed ($m=d_{\text{model}}$) and the graph is a permutation, a single head suffices. However, compression ($m>d_{\text{model}}$) forces relations into overlapping subspaces, creating interference that a single large head cannot disentangle. Our analysis shows that allocating a fixed $D_{K}$ across many small heads mitigates this interference, increasing the number of recoverable relations. Controlled single-layer experiments mirror the theory, revealing a sharp performance threshold that matches the predicted capacity scaling and confirms the benefit of distributing $D_{K}$ across multiple heads.
Altogether, these results provide a concrete scaling law for self-attention capacity and a principled design rule for allocating key-query budget across heads.
1 Introduction
The transformer architecture, and its self-attention mechanism in particular, has revolutionized fields from natural language processing to computer vision [60]. At its core, self-attention computes a similarity-weighted pattern of pairwise relationships among items in a context: queries match keys; the resulting scores route information via values [53, 36, 72, 61, 16, 57]. We ask a basic question: for a fixed self-attention mechanism size, how many target relationships can a single attention layer represent and reliably recover? We call this the layer’s capacity. Capacity is a foundational property for several reasons. (i) It imposes a hard ceiling on relational computation: beyond a threshold, no training procedure or dataset can make a layer uniformly recover all relations, much like rank bounds in linear models. (ii) It complements mechanistic work that isolates specific attention circuits in trained transformers [9, 63, 46, 33], by asking how many independent circuits can coexist. (iii) It provides an actionable resource scaling law, by describing how relationship capacity grows with increasing attention mechanism budget.
We also demonstrate that capacity impacts when increasing the number of heads is useful. Multiple heads are often conceived as a way for a source concept to attend to multiple different targets [60], but looking at self-attention through the lens of capacity demonstrates that multiple heads are beneficial even in the simple case where each concept attends to only a single target. Specifically, when compressed embeddings are used, many relations must be stored in overlapping subspaces; distributing the self-attention budget across many small heads reduces interference and increases the number of relations that can be cleanly separated—consistent with both pruning/specialization studies and expressivity results for attention [64, 45, 10].
One might hope to answer capacity empirically by probing large trained models. In practice, this is ill‑posed. Modern transformers superpose many relationships in shared subspaces; heads are polyfunctional and context‑dependent, so the number of “active” relations is not directly observable. Moreover, attention weights need not align with causal importance [28], and even sophisticated circuit‑tracing pipelines currently miss parts of the QK computation that determine where a head attends [33]. Beyond these methodological issues, superposition makes enumeration intrinsically hard: models can store more features than basis directions, packing multiple concepts into overlapping subspaces [18, 7]. As a result, interpretability work thus far has not revealed how many relationships can be supported by a fixed attention budget.
We therefore introduce a frameworkRelational Graph Recognition (RGR)—and analyze an idealized self‑attention model for solving RGR. The framework allows us to explicitly control both the structure and the number of attention relationships by casting self‑attention as recovering edges of a relational graph among $m$ items, while the model preserves the computational constraints and symmetries of attention. This allows predictions through principled analysis as well as controlled simulations that directly test those predictions. Our abstraction isolates the key–query computation that determines where a head attends, separating it from the OV pathway that determines what is written—a split made explicit in recent mechanistic analyses of attention heads [33]. As a result, our attention budget is defined in terms of the total key dimension $D_{K}=h\,d_{k}$, where $h$ is the number of heads and $d_{k}$ the per‑head key (and query) width.
Problem Formulation: Relational Graph Recognition (RGR)
To make “relationships” precise, we cast the core task of self-attention as a graph recovery problem.
Task.
Let $G=(V,E)$ be a directed graph on $m=\lvert V\rvert$ items with $m^{\prime}=\lvert E\rvert$ edges. A context is an ordered tuple $\mathcal{C}=(v_{i_{1}},\dots,v_{i_{\ell}})$ with $1\leq\ell\leq m$. The Relational Graph Recognition (RGR) mapping takes $(G,\mathcal{C})$ and, for each $v\in\mathcal{C}$, returns its in-context neighbors $N_{G}(v;\mathcal{C})=\{v^{\prime}\in\mathcal{C}:(v,v^{\prime})\in E\}.$ Given a graph $G$, we wish to find a parameterization $\Theta(G)$ such that the mapping defined by $\Theta(G)$ correctly produces $N_{G}(v;\mathcal{C})$ for every context $\mathcal{C}$ and every $v\in\mathcal{C}$.
Capacity question.
Fix an input embedding dimension $d_{\text{model}}$. For a graph family $\mathcal{G}_{m,m^{\prime}}=\{\,G:\lvert V\rvert=m,\ \lvert E\rvert=m^{\prime}\},$ we ask for the minimal total key dimension $D_{K}\;=\;h\,d_{k}$ (number of heads times per-head key/query width) such that the self-attention model in Section 3 can realize the RGR mapping for all $G\in\mathcal{G}_{m,m^{\prime}}$ and all contexts $\mathcal{C}$. We refer to this minimal $D_{K}$ as the capacity required by $\mathcal{G}_{m,m^{\prime}}$ at embedding dimension $d_{\text{model}}$.
Why this abstraction.
RGR isolates the key–query channel that determines where attention goes, while preserving the permutation symmetries and parameter sharing of self-attention [^1] We also describe below how positional embeddings can be incorporated into the model.. It lets us dial graph complexity $(m,m^{\prime})$ and the budget $D_{K}$ independently, enabling the information-theoretic bounds, constructive designs, and targeted experiments reported below. Exact mechanics (score computation, aggregation across heads, and omitted components such as softmax and values) are specified in Section 3.
Summary of Results
Our analysis yields both fundamental limits and constructive proofs of capability for self-attention as a relational reasoner, and our experiments validate the predictions in an idealized setting that mirrors the theoretical model. The main contributions are:
Formal model and budget.
We cast “where to attend” as Relational Graph Recognition (RGR) and analyze an idealized attention layer that preserves attention symmetries while isolating the key–query computation. The complexity measure is the total key dimension $D_{K}=h\,d_{k}$ (Sec. 3).
Information‑theoretic lower bound.
We show that recovering $m^{\prime}$-edge graphs on $m$ items with fixed margin requires
$D_{K}=\Omega\!\Big(\tfrac{m^{\prime}}{d_{\text{model}}}\log\tfrac{m^{2}}{m^{\prime}}\Big),$
independent of parameter precision and for any context length $\ell\geq 2$ (Sec. 7). This formalizes the intuition that the key-query matrices need to express sufficient information to describe the underlying relational graph. It also shows that representing more relationships demands greater key–query capacity, and that a smaller embedding dimension requires a larger total key dimension.
Asymptotically Optimal Constructions.
We provide explicit attention-based algorithms for RGR within the idealized model (Section 4). We achieve
$D_{K}\;=\;O\!\Big(\tfrac{(m^{\prime}+\Delta)}{d_{\text{model}}}\log m^{\prime}\Big),$
where $\Delta$ is the maximum degree of the graph $G$. Under a mild condition around balanced degrees ($(\Delta/d_{\text{avg}}\leq m/d_{\text{model}})$, this closes the gap between upper and lower bounds for all but very dense graphs, which have a gap factor of $\log m$. These constructions assume random Gaussian unit norm embeddings and also extend to any embedding satisfying a superposition hypothesis style near-orthogonality condition. These constructions also surface the core computational principles of self‑attention and serve as concrete, testable hypotheses about the internal mechanisms transformers might learn.
Capacity-Based Rationale for Multi-Head Attention.
Our analysis shows that multiple heads are beneficial even when each source has a single correct target (Section 5): when using compressed embeddings $d_{\text{model}}\ll m$, the signals for different relationships are superposed and multiple heads mitigate interference by specializing on disjoint subsets of relationships, thereby allowing the per-head dimension to be small. This provides a principled capacity-centric justification for multi-head attention as a method to reduce noise.
Empirical Validation of Capacity and Head Scaling.
In controlled single-layer experiments that mirror the idealized model, performance exhibits a sharp transition as $D_{K}$ increases (Section 6). The minimal $D_{K}$ needed to reach high accuracy (micro-F1 $\gtrsim 0.99$) grows rapidly with $m$ and shrinks with $d_{\text{model}}$, consistent with the theory. We recover the predicted scaling:
$\boxed{D_{K}^{\star}\;\approx\;C\,\frac{m\log m}{d_{\text{model}}}}$
with a single global threshold at test time. A one‑parameter fit to all data yields $C=1.19$ with $R^{2}=0.944$. Discarding three data points that the theory predicts to perform slightly worse yields $C=0.966$ with $R^{2}=0.992$. We also find a pronounced multi‑head advantage even for permutation graphs: the smallest passing models use several heads while keeping per‑head width small and the optimal head count scales linearly with $\frac{m}{d_{\text{model}}}$, as predicted by the theory. We also see that capacity is largely insensitive to context length over $\ell\!\in\!\{16,32\}$.
Together, these findings yield a quantitative scaling law for capacity as a function of $D_{K}$ and $d_{\text{model}}$, and reveal a principled multi‑head advantage even when each source has only a single correct target—clarifying when and how to allocate key–query budget across heads. The close alignment between our constructive bounds and empirical thresholds provides a concrete, falsifiable foundation for the computational principles that enable self-attention.
2 Related Work
Given the breadth of prior work on attention, we defer an extended survey to Appendix D, covering expressivity, language‑theoretic limits, connectivity, memorization, superposition, interpretability, and graph‑structured models.
Closest to our focus are works on memorization capacity [58, 35, 32], including analyses of memorization in attention modules [42]. While aligned in spirit, the problem formulations are different: memorization typically maps each context to a single output token/label, whereas our RGR setting asks for the recovery of in‑context neighbors for every context from a set of possible tokens. Reductions between the two would require memorization handling a combinatorial number of contexts (polynomial in $m$ for fixed $\ell$, and exponential when $\ell$ scales with $m$), and we are not aware of efficient reductions that preserve guarantees in either direction. Accordingly, bounds in one setting do not directly imply bounds in the other. Not surprisingly, capacity results for memorization provide different scaling laws then ours. For example they often include explicit dependence on sequence length $\ell$ [42, 35, 32], whereas in our analysis and experiments, length plays a limited role. Our abstraction isolates the key–query addressing step—“where to attend”—which mechanistic analyses identify as central to head routing [33]. In this sense, RGR complements parameter‑centric memorization settings that emphasize “what to output”: we target the capacity required to select the correct neighbors across contexts.
Beyond memorization, prior theory characterizes what Transformers can compute with sufficient resources—universality/approximation [70], fine‑grained attention‑matrix expressivity [39], and structural bottlenecks such as per‑head low rank and rank collapse without mixing [5, 15]. Language‑theoretic and composition results map limitations at fixed budgets [24, 47]; orthogonally, restricting connectivity rather than dimensions shows universality of $O(\ell)$‑sparse patterns and principled pruning of dense ones [71, 67]; and algorithmic views analyze in‑context procedures [38]. Mechanistic studies find head specialization and prune‑ability [9, 45], while memory‑centric views link attention and FFNs to associative/key–value memories [50, 22]. See Appendix D for more details.
3 Modeling the Self-Attention Mechanism
We model the key–query (QK) channel of a single self‑attention layer for RGR, retaining permutation symmetry and parameter sharing while omitting components that do not affect the binary edge decision (softmax, $1/\sqrt{d_{k}}$ scaling, and values; see Appendix. C). The input is an ordered context of distinct vertices of length $\ell\!\leq\!m$, given as $\mathcal{C}=(v_{i_{1}},\dots,v_{i_{\ell}})$, one each to $\ell$ attention units. Each $v\in V$ is described using a unique embedding $\mathbf{x}_{v}\in\mathbb{R}^{d_{\text{model}}}$. Positional information is not explicitly modeled; if needed, positions can be incorporated by treating $(\text{token},\text{position})$ as distinct vertices.
Single head.
Each attention unit with a single head uses the same shared projection matrices $W_{Q},W_{K}\in\mathbb{R}^{d_{\text{model}}\times d_{k}}$. For each $v_{i_{p}}\!\in\!\mathcal{C}$,
$\mathbf{q}_{i_{p}}=\mathbf{x}_{i_{p}}W_{Q},\qquad\mathbf{k}_{i_{p}}=\mathbf{x}_{i_{p}}W_{K}.$
The unnormalized score from source $v_{i_{p}}$ to target $v_{i_{q}}$ is
$S_{pq}=\mathbf{q}_{i_{p}}\cdot\mathbf{k}_{i_{q}}^{\top}.$
We declare an edge $(v_{i_{p}},v_{i_{q}})$ present iff $S_{pq}>\tau$ for a global threshold $\tau$. Only pairs inside $\mathcal{C}$ are tested.
Multi‑head extension.
With $h$ heads, each head $k$ has $(W_{Q}^{(k)},W_{K}^{(k)})\in\mathbb{R}^{d_{\text{model}}\times d_{k}}$ and produces $S_{pq}^{(k)}$. We aggregate per pair by
$S_{pq}^{\max}=\max_{k\in\{1,\dots,h\}}S_{pq}^{(k)},\quad\text{and decide}\quad(v_{i_{p}},v_{i_{q}})\in E\ \Leftrightarrow\ S_{pq}^{\max}>\tau.$
This “OR‑of‑heads” view matches the common specialization picture. Also, replacing $\max$ by log‑sum‑exp yields an equivalent classifier after a global threshold shift; see App. C.
Algorithmic objective and budget.
A construction for RGR maps a graph $G=(V,E)$ to weights $\{(W_{Q}^{(k)},W_{K}^{(k)})\}_{k=1}^{h}$ and a threshold $\tau$ that realize the correct edge decisions for all contexts $\mathcal{C}$, regardless of length $\ell$. We measure complexity by the total key dimension
$D_{K}\;=\;h\,d_{k},$
since QK is implemented as two batched multiplications with concatenated weights
$W_{Q}^{\text{cat}}=[W_{Q}^{(1)}|\cdots|W_{Q}^{(h)}]\in\mathbb{R}^{d_{\text{model}}\times(hd_{k})}\text{ and }W_{K}^{\text{cat}}=[W_{K}^{(1)}|\cdots|W_{K}^{(h)}]\in\mathbb{R}^{d_{\text{model}}\times(hd_{k})};$
the dominant cost and parameter footprint scale with $h\,d_{k}$ rather than $h$ or $d_{k}$ alone (see Appendix C). [^2] All statements remain (up to a factor of two) if one counts $D_{Q}+D_{K}$. Our goal is to minimize $D_{K}$ over a graph family $\mathcal{G}_{m,m^{\prime}}$ for a given $d_{\text{model}}$.
In Appendix C, we provide further justification for many of our model choices. Specifically, we address the use of thresholding instead of scaling/softmax, our head aggregation rule, as well as the lack of a value pathway.
4 Explicit Constructions for Relational Graph Recognition
In this section, we provide explicit constructions for the attention weights that solve RGR within the idealized model of Section 3. These constructions establish an upper bound on the $D_{K}$ required to solve RGR, providing a concrete measure of the self-attention mechanism’s capacity for this task. Each design is proven to perform with high probability (w.h.p.), meaning at least $1-m^{-\gamma}$ for some constant $\gamma>2$. We begin with a brief warm‑up sketch—recognizing permutation graphs with one‑hot embeddings—then move to compressive embeddings. We then generalize to arbitrary graphs. Throughout, $i$ indexes the source and $j$ a candidate target in $E$; keys are tied to targets and queries are tied to sources.
Construction I: Permutation Graphs with One‑Hot Embeddings (Warm-up/sketch).
We consider a permutation graph $G$ on $m$ items with edges $E=\{(v_{i},v_{\pi(i)})\}$ and one‑hot node encodings $\mathbf{x}_{i}=\mathbf{e}_{i}\in\mathbb{R}^{m}$ (so $d_{\text{model}}=m$). A single head ($h=1$) suffices. Assign each target $v_{j}$ a random binary signature as its key $\mathbf{k}_{j}$ by drawing $W_{K}\in\{0,1\}^{m\times d_{k}}$ with i.i.d. $\mathrm{Bernoulli}(p)$ entries (e.g., $p=1/4$) and setting $\mathbf{k}_{j}=\mathbf{e}_{j}W_{K}$. For each source $v_{i}$, set the query to the signature of its target: $\mathbf{q}_{i}=\mathbf{k}_{\pi(i)}$, i.e., the ii‑th row of $W_{Q}$ equals the $\pi(i)$‑th row of $W_{K}$. With $d_{k}=C\log m$ and threshold $\tau=\tfrac{p+p^{2}}{2}\,d_{k}$, one has
$S_{i,\pi(i)}=\mathbf{k}_{\pi(i)}\!\cdot\!\mathbf{k}_{\pi(i)}\sim\mathrm{Binomial}(d_{k},p),\qquad S_{ij}=\mathbf{k}_{\pi(i)}\!\cdot\!\mathbf{k}_{j}\sim\mathrm{Binomial}(d_{k},p^{2})\;\ (j\!\neq\!\pi(i)).$
Chernoff bounds and a union bound over all $(i,j)$ yield simultaneous separation $S_{i,\pi(i)}>\tau>S_{ij}$ w.h.p. when $d_{k}=\Theta(\log m)$, so a single head recovers all edges. This matches the $\Omega(\log m)$ lower bound (Section 7) and the same argument extends to softmax scoring (Appendix C). Full details—algorithm and proof—are deferred to Appendix B as a “warm‑up” proof; Construction II below generalizes this scheme to compressive embeddings, but first we point out that a separation of the the type described is sufficient to handle contexts of any length.
Lemma 4.1 (Context‑robustness).
Fix parameters $\{(W_{Q}^{(k)},W_{K}^{(k)})\}_{k=1}^{h}$ and a threshold $\tau$. Let the aggregated score be $S_{ij}^{\max}:=\max_{k}S_{ij}^{(k)}$. If, simultaneously for all $i\in V$,
$S_{i,\pi(i)}^{\max}>\tau\quad\text{and}\quad S_{ij}^{\max}<\tau\ \ \text{for all }j\neq\pi(i),$
then for every context $\mathcal{C}\subseteq V$ and every source $i\in\mathcal{C}$: (i) if $\pi(i)\in\mathcal{C}$ then $S_{i,\pi(i)}^{\max}>\tau$ and $S_{ij}^{\max}<\tau$ for all $j\in\mathcal{C}\!\setminus\!\{\pi(i)\}$; (ii) if $\pi(i)\notin\mathcal{C}$ then $S_{ij}^{\max}<\tau$ for all $j\in\mathcal{C}$. Hence the same parameters recognize $E|_{\mathcal{C}}$ for every $\mathcal{C}$ and every length $\ell$.
Proof.
Restricting from $V$ to $\mathcal{C}$ only removes candidate targets. If $\pi(i)\in\mathcal{C}$, the inequalities in $(\star)$ remain true after removing all $j\notin\mathcal{C}$. If $\pi(i)\notin\mathcal{C}$, every remaining $j\in\mathcal{C}$ satisfies $j\neq\pi(i)$, so $S_{ij}^{\max}<\tau$ by $(\star)$. ∎
Construction II: Permutations Under Compressive Embeddings
We now extend the permutation case to the compressive regime $d_{\text{model}}\ll m$ under a Gaussian unit‑norm embedding. [^3] Random spherical codes are nearly orthogonal—inner products concentrate tightly around 0—which lets us obtain clean dot‑product thresholds and $O(\log m)$ scaling in our separation arguments. This cosine‑geometry is also standard and effective in practice: many systems explicitly constrain features to a hypersphere (e.g., NormFace and ArcFace in face recognition; Spherical Text Embedding in NLP; spherical objectives in metric learning). See [62, 13] for concentration/JL background and [66, 14, 44, 73] for representative uses of unit‑sphere embeddings. Each item $v_{i}$ is embedded as a fixed vector $\mathbf{x}_{i}\in\mathbb{R}^{d_{\text{model}}}$ drawn i.i.d. as $\tilde{\mathbf{x}}_{i}\sim\mathcal{N}(0,I/d_{\text{model}})$ and then $L_{2}$‑normalized, i.e., $\mathbf{x}_{i}=\tilde{\mathbf{x}}_{i}/\|\tilde{\mathbf{x}}_{i}\|_{2}$. Write $X\in\mathbb{R}^{m\times d_{\text{model}}}$ for the matrix with ii‑th row $\mathbf{x}_{i}^{\top}$. Given such an embedding and permutation $\pi$, our goal is to construct attention parameters that recognize $G$.
Multi‑Head Algorithmic Construction
The fundamental challenge with embeddings is that the input $\mathbf{x}_{i}$ is a superposed representation of the node’s identity. Our construction first approximately inverts the embedding process, projecting the $d_{\text{model}}$-dimensional vector $\mathbf{x}_{i}$ back into the $m$-dimensional one-hot space using the transpose of the embedding matrix. We then apply the logic from the one-hot case. However, doing this with a single head yields too much noise due to the inversion being only approximate. We mitigate this noise by using multiple attention heads, where each is responsible for recognizing the outgoing edges from a disjoint subset of sources. This results in smaller individual heads, and thus less noise.
Algorithm 1 Construction for Permutation Graphs with Compressive Embeddings
1:Input: Permutation graph $G=(V,E)$ with $\pi:V\to V$; embedding matrix $X\in\mathbb{R}^{m\times d_{\text{model}}}$.
2:Parameters: Number of heads $h=\frac{m}{d_{\text{model}}}$; per‑head key/query dimension $d_{k}=C\log m$ for a sufficiently large absolute constant $C$. Set Threshold: $\tau=\frac{1}{2}d_{k}$.
3:Partition sources and targets: Split $V$ into $h$ disjoint blocks $V_{1},\dots,V_{h}$ of size $|V_{k}|=d_{\text{model}}$. For each head $k$, define its target set $T_{k}\;:=\;\{\pi(s):s\in V_{k}\},$ which is a permutation of $V_{k}$. Head $k$ is responsible for sources in $V_{k}$ and targets in $T_{k}$ (a permutation of $V_{k}$).
4:Random signatures: Draw $W_{\mathrm{sig}}\in\{\pm 1\}^{m\times d_{k}}$ with i.i.d. Rademacher entries; let $\mathbf{w}_{j}$ be its jj‑th row.
5:Ideal one‑hot‑space templates (for each head $k$):
6:   $\textbf{Query Matrix:}\;W^{\prime}_{Q,(k)}\in\mathbb{R}^{m\times d_{k}}$ with row $i$ equal to $\mathbf{w}_{\pi(i)}$ if $i\in V_{k}$, and $\mathbf{0}$ otherwise.
7:   $\textbf{Key Matrix:}\;W^{\prime}_{K,(k)}\in\mathbb{R}^{m\times d_{k}}$ with row $j$ equal to $\mathbf{w}_{j}$ if $j\in T_{k}$, and $\mathbf{0}$ otherwise.
8:Project back to model space (approximate de‑embedding):
$W^{(k)}_{Q}\;=\;X^{\top}W^{\prime}_{Q,(k)},\qquad W^{(k)}_{K}\;=\;X^{\top}W^{\prime}_{K,(k)}.$
Theorem 4.2 (Multi‑head recognition under Gaussian unit‑norm embeddings).
Assume the setup above with $h=\frac{m}{d_{\mathrm{model}}}$ heads, per‑head dimension $d_{k}=C\log m$ for a sufficiently large absolute constant $C$, and threshold $\tau=\tfrac{1}{2}d_{k}$. If $d_{\mathrm{model}}\geq c_{0}\log m$ for a sufficiently large absolute constant $c_{0}$, then with probability at least $1-m^{-3}$ over the draw of $(X,\text{signatures})$,
$\forall i\in V\;\exists\,k\in[h]\text{ with }i\in V_{k}:\quad S^{(k)}_{i,\pi(i)}>\tau\quad\text{and}\quad S^{(k)}_{ij}<\tau\ \ \forall j\neq\pi(i).$
Consequently, max‑pooling over heads correctly recognizes all edges and $D_{K}=h\,d_{k}=\Theta\!\Big(\frac{m\log m}{d_{\mathrm{model}}}\Big).$
Proof sketch (intuition).
Write $G=XX^{\top}$ and note $G\approx I$ w.h.p. when $d_{\text{model}}\gtrsim\log m$. For a source $i\in V_{k}$, the de‑embedded vector $\mathbf{u}_{i}:=\mathbf{x}_{i}X^{\top}$ equals $\mathbf{e}_{i}+\boldsymbol{\delta}_{i}$ where the leakage $\boldsymbol{\delta}_{i}$ has small $\ell_{2}$ mass on the head‑local blocks $V_{k}$ and $T_{k}$. Using Rademacher signatures, the score decomposes as
$S^{(k)}_{ij}=\underbrace{\mathbf{w}_{\pi(i)}\!\cdot\!\mathbf{w}_{j}\cdot\mathbb{I}(j\in T_{k})}_{\text{Signal}}\ +\ \underbrace{\text{terms linear/quadratic in }\boldsymbol{\delta}_{i},\boldsymbol{\delta}_{j}}_{\text{Noise}}.$
At the true edge $j=\pi(i)$, Signal equals $d_{k}$ exactly; at non‑edges it is 0 (if $j\notin T_{k}$) or a sub‑Gaussian fluctuation $O(\sqrt{d_{k}\log m})$ (if $j\in T_{k}\setminus\{\pi(i)\}$). Concentration results for random spherical codes implies (i) per‑head leakage energy $\|\boldsymbol{\delta}_{i,S}\|_{2}^{2}\lesssim 1$ for $S\in\{V_{k}\!\setminus\!\{i\},T_{k}\}$ and (ii) small cross‑terms $\langle\boldsymbol{\delta}_{i,\pi^{-1}(T_{k})},\boldsymbol{\delta}_{j,T_{k}}\rangle\lesssim\sqrt{\frac{\log m}{d_{\text{model}}}}$. Thus each Noise component concentrates to $O(\sqrt{d_{k}\log m})+o(d_{k})$; choosing $d_{k}=C\log m$ and $d_{\text{model}}\geq c_{0}\log m$ makes Noise $\tfrac{1}{4}d_{k}$, while Signal at the true edge is $d_{k}$. Setting $\tau=\tfrac{1}{2}d_{k}$ yields $S^{(k)}_{i,\pi(i)}>\tau>S^{(k)}_{ij}$ uniformly, and a union bound over $(i,j,k)$ gives the stated probability. A full proof with explicit constants and concentration lemmas appears in Appendix B.
Consequences.
The total key budget satisfies $D_{K}=O\!\big(\frac{m\log m}{d_{\text{model}}}\big)$, which matches our lower bound up to constants. Moreover, the non‑edge bounds hold head‑wise, so for any $j\neq\pi(i)$ we have $S_{ij}^{(k)}<\tau$ for all $k$ and hence $S_{ij}^{\max}<\tau$, while $S_{i,\pi(i)}^{\max}>\tau$ at the true target. By Lemma 4.1, the same parameters correctly recognize $E|_{\mathcal{C}}$ for every context $\mathcal{C}$ and for every context length. (As in Construction I, the thresholding analysis translates to softmax; see Section C.)
Construction III: More General Embeddings (Summary).
In Appendix B, we abstract the geometric assumptions used in Construction II into a reusable, block‑level condition on the embedding $X$: restricted self‑incoherence (Def. B.4), which requires the existance of a block size parameter $B$ such that (a) after rescaling by a global factor $\mu$, the Gram matrix $XX^{\top}$ has diagonals close to $1$; (b) when a row of $XX^{\top}$ is restricted to any subset of at most $B$ indices, the off‑diagonal “leakage” energy is small (so projecting $\mathbf{x}_{i}$ back with $X_{\mathrm{inv}}:=\tfrac{1}{\mu}X^{\top}$—a scaled transpose acting as an approximate inverse—yields a near one‑hot vector on any BB‑sized block); and (c) the leakage patterns of two different rows have small inner product on any such block (small “cross‑leakage”). Given these conditions, we synthesize reference one‑hot weight matrices for keys/queries on each block of size $B$, and realize them in model space by multiplying with $X_{\mathrm{inv}}$. With sparse Bernoulli or Rademacher signatures of width $d_{k}=\Theta(\log m)$ and a fixed dot‑product threshold, each head cleanly separates the true target within its block; max‑pooling over $h\asymp m/B$ heads then recognizes all edges w.h.p., using a total key budget $D_{K}=\Theta\!\big((m/B)\log m\big)$.
Informally, our assumptions are superposition‑style: many items are encoded in a shared low‑dimensional space so that a single linear decoder approximately recovers each one‑hot identity while keeping interference within any block of at most $B$ items uniformly small. Technically, this is a block‑wise low‑coherence (JL/RIP‑style) requirement on $XX^{\top}$, not a general superposition claim. This framework recovers Construction II as a special case: for Gaussian unit–norm embeddings (Cor. B.6) one may take $B=\Theta(d_{\text{model}})$, giving $D_{K}=\Theta\!\big(m\log m/d_{\text{model}}\big)$; for sparse random binary compressive embeddings with $p_{B}=\Theta(\log m/d_{\text{model}})$ (Cor. B.7), one can take $B=\Theta(d_{\text{model}}/\log m)$, giving $D_{K}=\Theta\!\big(m\log^{2}m/d_{\text{model}}\big)$. Full algorithms, constants, and proofs appear in Appendix B.
Construction IV: General Graphs (Summary).
We also extend the compressive permutation scheme (Construction II) to any directed graph $G=(V,E)$ with $|V|=m$ and $|E|=m^{\prime}$. First, we pack $E$ into disjoint matchings of size at most $d_{\text{model}}$ by edge‑coloring the bipartite incidence graph (Kőnig) and batching: this yields $H=\Big\lceil\tfrac{m^{\prime}}{d_{\text{model}}}\Big\rceil+\Delta$ heads, where $\Delta=\max\{\Delta_{\mathrm{out}},\Delta_{\mathrm{in}}\}$. Head $k$ operates on one matching $M_{k}$ (a partial permutation $\pi_{k}:V_{k}\!\to\!T_{k}$) and reuses the Construction II machinery: shared Rademacher signatures, approximate de‑embedding via $X^{\top}$ under Gaussian unit‑norm embeddings, per‑head dimension $d_{k}=C\log m$, and a global threshold $\tau=\tfrac{1}{2}d_{k}$. Exactly one head contains each true edge, while non‑edges lie below $\tau$ in every head; with $d_{\text{model}}\!\geq\!c_{0}\log m$, concentration and a union bound give uniform separation $S^{(k)}_{i,\pi_{k}(i)}>\tau>S^{(k)}_{ij}$ w.h.p., so max‑pooling recovers $E$ and inherits context‑robustness (Lemma 4.1). The total key budget is
$D_{K}\;=\;O\!\left(\frac{m^{\prime}\log m}{d_{\text{model}}}\;+\;\Delta\log m\right).$
Under the mild skew condition $\Delta\leq\frac{m^{\prime}}{d_{\text{model}}}$ (equivalently $\Delta/d_{\mathrm{avg}}\leq m/d_{\text{model}}$ with $d_{\mathrm{avg}}=m^{\prime}/m$), we can take $H=\lceil m^{\prime}/d_{\text{model}}\rceil$ and obtain the tighter $D_{K}=\Theta\!\left(\frac{m^{\prime}\log m^{\prime}}{d_{\text{model}}}\right),$ matching the information‑theoretic lower bound up to constants for all but the densest graphs. Full algorithms, constants, and proofs are deferred to Appendix B.
5 The power of multiple heads
[…]
$N_{3}(B)\;\asymp\;\frac{B}{d_{\text{model}}}\,\sqrt{d_{k}\log m}.$ (1)
To guarantee (w.h.p.) a fixed margin between the true target and all non‑targets, it must be that, for constants $c_{1},c_{2}>0$,
$N_{3}(B)\;\leq\;c_{1}d_{k}\quad\Longrightarrow\quad d_{k}\;\geq\;c_{2}\,\frac{B^{2}}{d_{\text{model}}^{2}}\,\log m.$ (2)
Single head with compression. If one head serves all items, then $B=m$. Applying (2) gives the requirement $d_{k}\;\geq\;c_{2}\,\frac{m^{2}}{d_{\text{model}}^{2}}\,\log m,$ and since $h=1$ here, the total key dimension is $D_{K}=d_{k}$.
Multiple heads with compression. Construction II partitions the items into $\frac{m}{d_{\text{model}}}$ heads with $B=d_{\text{model}}$ per head. Plugging $B=d_{\text{model}}$ into (1) yields $N_{3}(B)=\Theta(\sqrt{d_{k}\log m})$. Taking $d_{k}=c_{3}\log m$ with $c_{3}$ larger than the constant in (1) ensures $N_{3}\leq c_{1}d_{k}$ w.h.p., and the total key dimension is $D_{K}\;=\;h\,d_{k}\;=\;O\!\Big(\frac{m}{d_{\text{model}}}\log m\Big).$
Consequence. As a result, in the compressive regime, if $m=\omega(d_{\text{model}})$, the single head requirement above implies asymptotically larger $D_{k}$ than the multihead construction. Or, equivalently, for a fixed $D_{k}$ budget, multiple heads can handle more edges (relationships) than a single head, even in a permutation graph. Multiple heads do not boost per‑head expressivity; they localize de‑embedding noise by reducing block size $B$, so that each head aggregates leakage over only fewer coordinates, bringing the noise to a manageable level. Note that this is not a lower bound for all conceivable single‑head designs, but it shows that within the de‑embedding to signature template we use, a single head cannot perform as well as multiple heads.
6 Experiments
We conduct experiments in an idealized self-attention setting, mirroring our theoretical model, to test several predictions. First, we compare the empirical minimum total key dimension, $\hat{D}_{K}^{\star}$, to the predicted theoretical scaling law of $\Theta\left(\frac{m\log m}{d_{\text{model}}}\right)$, noting that optimization may fail to find a solution matching the theoretical constructive bound. And second, we test predictions regarding head count: whether a multi-head advantage appears in permutation graphs and how the empirically optimal number of heads tracks the theory.
Experimental implementation
We empirically instantiate the idealized attention layer of our framework with two learned projections $W_{Q},W_{K}\in\mathbb{R}^{d_{\text{model}}\times D_{K}}$ partitioned into $h$ heads ($d_{k}=D_{K}/h$). For a context matrix $X_{\mathcal{C}}$, head $k$ computes $S^{(k)}=Q^{(k)}(K^{(k)})^{\top}$ with $Q^{(k)}=X_{\mathcal{C}}W_{Q}^{(k)}$ and $K^{(k)}=X_{\mathcal{C}}W_{K}^{(k)}$; scores are combined by an elementwise max $S_{\max}=\max_{k}S^{(k)}$, and we predict an edge $(p\!\to\!q)$ iff $S_{\max}(p,q)>\tau$ for a single learned global threshold $\tau$. There is no $1/\sqrt{d_{k}}$ scaling, softmax, or value pathway, so capacity is purely key–query. Tasks are permutation graphs on $m$ items (one out/in-edge per node). Node embeddings $x_{i}\!\sim\!\mathcal{N}(0,I/d_{\text{model}})$ are $L_{2}$‑normalized and frozen, making $D_{K}$ the sole capacity knob. Contexts of length $\ell$ (default $\ell{=}16$) are sampled with target‑in‑context rate $\rho{=}0.5$.
We train $W_{Q},W_{K},\tau$ with AdamW (lr $10^{-3}$, weight decay 0) using a weighted logistic loss over all ordered pairs within a context (positive weight $\ell{-}1$; logit sharpness $\alpha{=}10$), one context per step. For each run, a single permutation $\pi$ and embedding matrix are fixed by seed; training contexts are drawn on‑the‑fly, with 500 validation and 2,000 held‑out test contexts from the same $(\ell,\rho)$ distribution. Early stopping checks validation micro‑F1 every 500 steps and halts after five consecutive checks above $0.995$. We report micro‑F1 on the fixed test set with the single learned $\tau$; the “minimum $D_{K}$” is the smallest $D_{K}$ achieving mean test micro‑F1 $\geq 0.99$ for at least one head count $h$. Full details appear in App. A.1.
6.1 Results and comparison to theory
We probe capacity on permutation graphs with $m\!\in\!\{64,128,256,512\}$ and $d_{\text{model}}\!\in\!\{16,32,64\}$. For each $(m,d_{\text{model}})$ we sweep head counts $h\!\in\!\{1,2,4,8,16,32,64\}$ and several total key sizes $D_{K}{=}h\,d_{k}$ (multiple $D_{K}$ per $h$). Each configuration is trained from $10$ seeds with the protocol described above (AdamW, fixed embeddings, single global threshold $\tau$). We evaluate average test micro–F1 on a fixed held‑out set and define the empirical threshold
$D_{K}^{\star}\;=\;\min\{\,D_{K}\;:\;\exists h\ \text{s.t. mean test micro‑F1}\geq 0.99\,\}.$
We denote by $h^{\star}$ a head count that attains $D_{K}^{\star}$. Full grids and per‑config step limits are in App. A.2.
To isolate sequence‑length effects at fixed embedding compression, we also traverse the diagonal $r\!\stackrel{{\scriptstyle\text{def}}}{{=}}\!m/d_{\text{model}}{=}8$ with $(m,d_{\text{model}})\in\{(128,16),(256,32),(512,64),(1024,128),(2048,256),(4096,512)\}$, using $3$ seeds for the largest points and increased budgets (App. A.2). Finally, to probe extreme compression we include a second r=32r{=}32 point $(1024,32)$ (in addition to $(512,16)$). We also repeat a subset of runs with a variable context length (App. A.3).
Qualitative phenomena.
We observe: (i) a sharp F1 transition in a narrow $D_{K}$ window (capacity threshold) across all $(m,d_{\text{model}},h)$ (Fig. 2 below and Fig. 6 in App. A.2); (ii) a pronounced multi‑head advantage for many $(m,d_{\text{model}})$, even though each query has a single target—splitting a fixed $D_{K}$ across more heads reduces interference from superposition (Fig. 2); and (iii) the optimal head count increases with compression $r{=}m/d_{\text{model}}$ (Fig. 2), while per‑head width at the threshold remains modest.
[…]
Figure 2: More heads are needed as $m$ grows and as $d_{\text{model}}$ shrinks. See App. A.2 for error bar description.
Empirical thresholds on the base grid.
The minimum $D_{K}^{\star}$ grows rapidly with $m$ and decreases rapidly with $d_{\text{model}}$; exact values appear in Figure 5 (in the Appendix). A single head often fails to reach $0.99$ F1 within the scanned $D_{K}$ (e.g., $(m,d_{\text{model}})\!\in\!\{(512,64),(256,32)\}$, Fig. 6), whereas several small heads pass at substantially smaller $D_{K}$.
Scaling laws
Plotting $D_{K}^{\star}$ against $\frac{m\log m}{d_{\text{model}}}$ yields a tight linear relation (Fig. 4):
$D_{K}^{\star}\;\approx\;\mathbf{1.19}\cdot\frac{m\log m}{d_{\text{model}}}\qquad(R^{2}=\mathbf{0.944}).$
We see small deviations when $d_{\text{model}}$ is too small relative to $\log m$; this is consistent with our theoretical results. Excluding the three $(d_{\text{model}}{=}16,\,m{>}64)$ points (above the line in Fig 4)—which violate the precondition $d_{\text{model}}\!\gtrsim\!c_{0}\log m$ used by our constructions—gives slope $0.966$ with $R^{2}{=}\mathbf{0.992}$. Thus, the empirical capacity closely matches the theoretical $\Theta\!\big(\tfrac{m\log m}{d_{\text{model}}}\big)$ rate. The head count that attains $D_{K}^{\star}$ scales approximately linearly with compression (Fig. 7 in the Appendix):
$h^{\star}\;\approx\;\mathbf{1.65}\,\frac{m}{d_{\text{model}}}\;-\;\mathbf{6.64}\qquad(R^{2}=\mathbf{0.824}).$
At $D_{K}^{\star}$, per‑head widths are small: $d_{k}^{\star}\!\in\![5,24]$ on the base grid (median $11$), indicating gains come primarily from adding heads rather than making each head wide (Table 1; App. A.2).
Figure 3: Comparison of $D_{K}^{\star}$ to theoretical scaling law. $x$-axis: scaling law prediction. $y$-axis: observed behavior. See App. A.2 for error bar description.
[…]
Fixed‑compression diagonal (r=8r{=}8).
Holding $r$ constant collapses the prediction to $D_{K}^{\star}\propto r\log m$, so the dependence on $m$ should be logarithmic. Along $(128,16)\!\to\!(4096,512)$ we observe roughly this behavior from $m{\geq}512$ onward (Fig. 4): $D_{K}^{\star}$ grows slowly while $m$ grows exponentially, matching the $\log m$ factor. The first two points are slightly conservative (smaller $d_{\text{model}}$) and align with the same $d_{\text{model}}\!\gtrsim\!\log m$ finite‑size effect. We also expect the optimal head count to be proportional to $r$; the observed results align well with this expectation.
Takeaways.
Empirical thresholds align closely with the $m\log m/d_{\text{model}}$ capacity law and expose a clear multi‑head advantage even for one‑target graphs. Discrepancies appear exactly where theory anticipates stronger superposition (small $d_{\text{model}}$ and very large $m$). Overall, allocating key–query budget across more heads with modest width is the efficient path to capacity in compressed embeddings. In Appendix A.3, provide evidence that these results are largely independent of context length. Specifically, thresholds were stable across $\ell\in\{16,32\}$ with a small shift only when testing at longer $\ell$ than used for training.
7 Lower Bound on Relational Graph Recognition
In this section, we provide our lower bound on any self‑attention mechanism that uniformly recovers every directed graph on $m$ items with exactly $m^{\prime}$ edges in our model from Section 3 under a fixed positive margin $\gamma$. The number of such graphs is $\binom{m(m-1)}{m^{\prime}}$; essentially what we show is that the QK parameters must carry (at least) the description length of the edge set, and thus the total key dimension
$D_{K}\;=\;\Omega\!\Big(\tfrac{\log\binom{m(m-1)}{m^{\prime}}}{d_{\text{model}}}\Big)=\Omega\Big(\tfrac{m^{\prime}\log(m^{2}/m^{\prime})}{d_{\text{model}}}\Big).$
The result is independent of parameter precision and applies to any context length $\ell\geq 2$.
We start with some preliminaries. We first point out that we can focus only on length‑2 contexts. Uniform correctness for RGR requires that, for every ordered pair $(u,v)\in V\times V$, the decision $(u,v)\in E$?” is the same in every context containing $u$ and $v$. In particular it must be correct in the length‑2 context $\mathcal{C}=(u,v)$. Hence any lower bound proved using only length‑2 contexts applies to the full problem. We next provide two structural reductions.
(i) Multi‑head to a single bilinear form. For a length‑2 context and any monotone per‑pair head aggregator (max, log‑sum‑exp, sum), replacing it by the sum only makes the model more permissive. Writing
$M\;=\;\sum_{k=1}^{h}W_{Q}^{(k)}W_{K}^{(k)\top}\ \in\ \mathbb{R}^{d_{\text{model}}\times d_{\text{model}}},$
the score on $(u,v)$ is the single bilinear form
$S(u,v)\;=\;\mathbf{x}_{u}^{\top}M\,\mathbf{x}_{v},$ (3)
and an edge is recognized iff $S(u,v)>\tau$ for a global threshold $\tau$. Since each summand $W_{Q}^{(k)}W_{K}^{(k)\top}$ has rank at most $d_{k}$,
$\mathrm{rank}(M)\ \leq\ \sum_{k=1}^{h}d_{k}\;=\;D_{K}.$ (4)
[…]
(ii) Edge‑set description length. Let $m=|V|$ and N=m(m1)N=m(m-1) be the number of ordered, loop‑free pairs. For any target size $m^{\prime}$, the family $\mathcal{G}_{m,m^{\prime}}$ of directed graphs with exactly $m^{\prime}$ edges has cardinality $\binom{N}{m^{\prime}}$; hence any procedure that can realize an arbitrary $E\subseteq[N]$ with $|E|=m^{\prime}$ must (implicitly) transmit at least
$L(m,m^{\prime})\;=\;\ln\binom{N}{m^{\prime}}$ (5)
[…]
Definition 7.1 (Constant‑margin recovery on length‑2 contexts).
A parameter choice $\big(\{W_{Q}^{(k)},W_{K}^{(k)}\}_{k=1}^{h},\tau\big)$ recovers a graph $G=(V,E)$ with margin $\gamma>0$ if, for every ordered pair $(u,v)$ with $u\neq v$,
$(u,v)\in E\Rightarrow\mathbf{x}_{u}^{\top}M\,\mathbf{x}_{v}\geq\tau+\gamma,\qquad(u,v)\notin E\Rightarrow\mathbf{x}_{u}^{\top}M\,\mathbf{x}_{v}\leq\tau-\gamma,$
where $M=\sum_{k}W_{Q}^{(k)}W_{K}^{(k)\top}$.
Because the decision rule is homogeneous in $(M,\tau)$, we fix scale by requiring
$M=UV^{\top},\quad\|U\|_{F}\leq 1,\ \|V\|_{F}\leq 1,\ \ |\tau|\leq 1.$ (6)
Under (6) and $\|\mathbf{x}_{u}\|_{2},\|\mathbf{x}_{v}\|_{2}\leq 1$, the score map is $1$-Lipschitz in Frobenius norm:
$\big|\mathbf{x}_{u}^{\top}UV^{\top}\mathbf{x}_{v}-\mathbf{x}_{u}^{\top}\tilde{U}\tilde{V}^{\top}\mathbf{x}_{v}\big|\ \leq\ \|U-\tilde{U}\|_{F}+\|V-\tilde{V}\|_{F},\qquad\text{and }\ |(\tau-\tilde{\tau})|\ \text{adds linearly.}$ (7)
Therefore any perturbation of $(U,V,\tau)$ of radius at most $\gamma/4$ preserves all pairwise signs, and hence the entire edge set, on length‑2 contexts.
Theorem 7.2 (Description‑length lower bound for QK).
Fix $m\in\mathbb{N}$ and $m^{\prime}\in\{0,\dots,m(m\!-\!1)\}$. Suppose an attention mechanism of the form (3)–(4), with item embeddings $\|\mathbf{x}_{v}\|_{2}\leq 1$, can recover every graph in $\mathcal{G}_{m,m^{\prime}}$ with margin $\gamma\in(0,1)$. Then there exists a constant $c(\gamma)>0$ such that
$d_{\text{model}}\;D_{K}\ \ \geq\ \ c(\gamma)\,\log\binom{m(m-1)}{m^{\prime}}\ -\ O(1).$ (8)
Equivalently,
$D_{K}\ \ =\ \ \Omega\!\Bigg(\frac{\log\binom{m(m-1)}{m^{\prime}}}{d_{\text{model}}}\Bigg).$ (9)
Proof.
Under the normalization (6) and Lipschitz property (7), the parameter set $\mathbb{B}=\{(U,V,\tau):\|U\|_{F},\|V\|_{F},|\tau|\leq 1\}$ admits an $\varepsilon$-net of radius $\varepsilon=\gamma/4$ of size at most
$N_{\mathrm{cov}}(\varepsilon)\ \leq\ \Big(\tfrac{C}{\varepsilon}\Big)^{\,2\,d_{\text{model}}D_{K}+1}\ =\ \Big(\tfrac{C^{\prime}}{\gamma}\Big)^{\,2\,d_{\text{model}}D_{K}+1}$
for absolute constants $C,C^{\prime}>0$. Each net point induces a unique labeling of the N=m(m1)N=m(m-1) ordered pairs by the margin, hence at most $N_{\mathrm{cov}}(\varepsilon)$ distinct edge sets can be realized. Since the mechanism must realize all $\binom{N}{m^{\prime}}$ edge sets of size $m^{\prime}$, we obtain $\binom{N}{m^{\prime}}\leq N_{\mathrm{cov}}(\varepsilon)$, which rearranges to (8)–(9). ∎
Equivalent finite‑precision statement.
If each real parameter in $\{W_{Q}^{(k)},W_{K}^{(k)},\tau\}$ has $b=\Theta(1)$ effective bits after normalization (e.g., due to quantization or stochastic rounding), then the parameter budget contains at most $B=b\,(2\,d_{\text{model}}D_{K}+1)$ bits and thus can realize at most $2^{B}$ distinct edge sets. Requiring $2^{B}\geq\binom{N}{m^{\prime}}$ gives the same conclusion as (9) with an explicit constant $1/(2b)$. Under the margin model above, one may take $b=\Theta(\log(1/\gamma))$.
Bounds for specific cases.
This demonstrates the following results:
  • Exactly $m^{\prime}=m$ edges (such as permutation graphs): $D_{K}=\Omega\Big(\frac{m^{\prime}\,\log m}{d_{\text{model}}}\Big).$
  • Dense regime with $m^{\prime}=\Theta(m^{2})$: $D_{K}=\Omega\Big(\frac{m^{\prime}}{d_{\text{model}}}\Big)$.
  • Sparse regime with $m^{\prime}=O(m^{2-\epsilon})$ for some positive constant $\epsilon$: $D_{K}=\Omega\Big(\frac{m^{\prime}\,\log m}{d_{\text{model}}}\Big)$.
8 Limitations
Theory. Our constructive upper bounds are tight only under a mild degree–skew assumption (e.g., $\Delta/d_{\text{avg}}\leq m/d_{\text{model}}$). When this condition is violated—or in very dense graphs—the present upper bounds do not match the information-theoretic lower bound. Moreover, our sufficiency results rely on geometric properties of the embeddings (near-orthogonality / restricted self-incoherence). While we verify these conditions for idealized random embeddings, we do not quantify how embeddings learned in trained models satisfy (or deviate from) these properties, nor the effect of such deviations on the constants in our bounds.
Experiments. Empirically we restrict to permutation graphs (out-degree =1=1) and do not test graphs with larger out-degree. We also evaluate a QK-only layer and thus do not experimentally validate modeling assumptions such as adding a value pathway or incorporating a softmax decision rule. Our search over $D_{K}$ is coarse because $D_{K}$ is tuned in multiples of the head count; there remains room to probe all cross-points in the scanned range and to explore substantially larger dimensions, but we did not do so due to compute limits. Finally, while capacity appeared largely insensitive to context length within $\ell\in\{16,32\}$, we did not study much longer contexts for the same reason.
Appendix A More details on our experiments
A.1 Experimental implementation Details
This appendix reproduces the full experimental protocol (model specification, context sampling procedure, loss, optimization, early stopping, and evaluation criteria) described in Section 6.
Our experiments instantiate the upper‑bound model from Section 3 as follows. The parameters are two learned projections $W_{Q},W_{K}\in\mathbb{R}^{d_{\text{model}}\times D_{K}}$ and a single global scalar threshold $\tau$. We conceptualize $W_{Q},W_{K}$ as $h$ head blocks of width $d_{k}=D_{K}/h$. Scoring is done for a context matrix $X_{\mathcal{C}}\in\mathbb{R}^{\ell\times d_{\text{model}}}$, where head $k$ produces $S^{(k)}=Q^{(k)}(K^{(k)})^{\top}\in\mathbb{R}^{\ell\times\ell}$ with $Q^{(k)}=X_{\mathcal{C}}W_{Q}^{(k)}$ and $K^{(k)}=X_{\mathcal{C}}W_{K}^{(k)}$. Scores are aggregated by elementwise max across heads: $S_{\max}=\max_{k}S^{(k)}$. At evaluation time we predict an edge $(p\!\to\!q)$ iff $S_{\max}(p,q)>\tau$. This matches the theoretical mechanism exactly: there is no $1/\sqrt{d_{k}}$ scaling, no softmax, and no value pathway—so capacity is purely key–query driven.
Our primary testbed is the family of permutation graphs $\{(V,E_{\pi})\}$ with $|V|=m$ and $E_{\pi}=\{(i,\pi(i)):\,i\in V\}$, where $\pi$ is a uniformly random permutation. This realizes the $m^{\prime}=m$ constructive case used in our upper bound and isolates the single‑target setting in which head specialization is most interpretable. Consistent with our constructions, each node $i\in V$ has a fixed embedding $x_{i}\in\mathbb{R}^{d_{\text{model}}}$ drawn i.i.d. from $\mathcal{N}(0,I/d_{\text{model}})$ and then $L_{2}$-normalized. Embeddings are frozen throughout training and evaluation. This both aligns with the random (nearly orthogonal) embedding assumption in our proofs and makes $D_{K}$ the sole capacity knob. An example is a context $\mathcal{C}$ of length $\ell$ (baseline $\ell=16$). To prevent degenerate class imbalance when $\pi(i)$ often falls outside $\mathcal{C}$, we enforce a target per‑context positive rate $\rho\in(0,1)$ as follows:
  1. 1.
    Sample a set $S\subseteq V$ of $\ell$ distinct nodes uniformly.
  2. 2.
    Sample $b\sim\mathrm{Binomial}(\ell,\rho)$ and choose $U\subseteq S$ with $|U|=b$.
  3. 3.
    For each $i\in U$, if $\pi(i)\notin S$ replace a random $j\in S\setminus\{i\}$ by $\pi(i)$, preserving $|S|=\ell$ and distinctness.
All of our experiments use $\rho=0.5$. This preserves the RGR semantics (positives remain exactly those $(i,\pi(i))$ that land in the same context) while reducing the time required to train.
For each experiment we sample one permutation $\pi$ and one embedding matrix $X$ using a fixed seed. We then generate a validation set of 500 contexts and a held‑out test set of 2,000 contexts with the same $(\ell,\rho)$ distribution. We then draw training contexts on the fly from the same generator (one context per optimization step).
We train $W_{Q},W_{K},\tau$ by minimizing a weighted logistic loss on all ordered pairs within a context:
$z_{pq}\;=\;\alpha\big(S_{\max}(p,q)-\tau\big),\qquad\mathcal{L}\;=\;\tfrac{1}{|\mathcal{C}|^{2}}\!\sum_{p,q}\!\Big[\underbrace{\mathrm{softplus}(-z_{pq})y_{pq}}_{\text{positive term}}\cdot\underbrace{\mathrm{pos_{weight}}}_{=\ell-1}+\mathrm{softplus}(z_{pq})(1-y_{pq})\Big],$
where $y_{pq}=1$ iff $(v_{i_{p}},v_{i_{q}})\in E$. The weighting $\mathrm{pos_{weight}}=\ell-1$ reflects that each source has at most one positive among $\ell$ candidates.
We use AdamW with learning rate $10^{-3}$ and weight decay 0. Parameters are initialized with $W_{Q},W_{K}\sim\mathcal{N}(0,1/\sqrt{d_{\text{model}}})$ and $\tau=0$. The logit sharpness is $\alpha=10$. We train for a number of steps with early stopping: every 500 steps we compute validation micro‑F1; if it exceeds $0.995$ for 5 consecutive checks, training halts. The number of steps increases with problem complexity. We use one context per step (contexts are small and independent), which keeps the implementation close to the theoretical algorithm and avoids artifacts from large mini‑batches.
All evaluation is conducted on the fixed held‑out test set of 2,000 contexts using the single learned threshold $\tau$ shared across all contexts. Our metric is Micro‑F1 over all ordered pairs across all test contexts. This directly measures correctness of binary edge recognition per the RGR objective. While the stopping rule uses validation F1 >0.995>0.995, the minimum $D_{K}$ we report below is extracted on the test set using a looser criterion: the smallest $D_{K}$ achieving mean micro–F1 $\geq 0.99$ for at least one $h$. We use $0.99$ to keep a margin from the stopping rule. All tables and statements about minimum $D_{K}$ are based on this $0.99$ test criterion.
A.2 Result details
We provide more detail on the results we found, additional details on the configurations used to find them, as well as the methodology we used for error bar determinination.
Methodology for Error Interval Construction
Display CIs for F1 curves.
Unless otherwise noted, error bars are 95% $t$-intervals across seeds: $\bar{F}_{1}\pm t_{0.975,\,n-1}\,s/\sqrt{n}$, where $n$ is the number of runs and $s$ their sample standard deviation. Intervals reflect training-run variability with a fixed test set.
Minimum key dimension $D_{K}^{\star}$.
The error interval for the minimum total key dimension, $D_{K}^{\star}$, is designed to reflect the uncertainty in the F1 score. For any given model configuration, we determine a central estimate along with an optimistic lower bound and a conservative upper bound, all based on a required F1 score of at least 0.99.
Let the mean F1 score from a set of trials be $\bar{F_{1}}$, with its corresponding 95% confidence interval being $F_{1,\text{low}},F_{1,\text{high}}]$. The three reported values for $D_{K}^{\star}$ are defined as follows:
  • Central Estimate: The primary value reported. It’s the minimum $D_{K}$ found for which the mean F1 score meets the performance threshold ($\bar{F_{1}}\geq 0.99$).
  • Conservative Upper Bound: This is the minimum $D_{K}$ for which the lower bound of the F1 confidence interval meets the threshold ($F_{1,\text{low}}\geq 0.99$). This stricter condition identifies the $D_{K}$ needed to be 95% confident that the true performance is sufficient.
  • Optimistic Lower Bound: This is the minimum $D_{K}$ for which the upper bound of the F1 confidence interval meets the threshold ($F_{1,\text{high}}\geq 0.99$). This looser condition identifies the $D_{K}$ for which it is merely plausible that the true performance is sufficient.
Optimal number of heads.
Let $h^{\star}$ be the head count achieving $D_{K}^{\star}$ (ties broken by larger $\bar{F}_{1}$). We form a candidate pool of head counts whose tested $D_{K}$ lies within $10\%$ of $D_{K}^{\star}$. Each candidate is compared to $h^{\star}$ using a paired two-sided $t$-test on per-seed F1; candidates with $p>0.05$ are labeled “not significantly different” and retained. [^4] We do not interpret $p>0.05$ as proof of equivalence; it only indicates insufficient evidence of a difference at $\alpha=0.05$. The reported interval spans the minimum and maximum head counts retained.
Further Details on Results
Figure 5: Minimum total key dimension $D_{K}^{\star}$. Upper right and lower left numbers represent confidence range; methodology described in the text.
Figure 5 lists $D_{K}^{\star}$, the minimum total key dimension, found for each configuration of $m$ and $d_{\text{model}}$ we tested. We use our minimum key dimension $D_{K}^{\star}$ intervals methodology, with the upper right corner being the upper bound and the lower left corner being the lower bound. The $d_{K}^{\star}$ (per head key size) used to achieve these $D_{K}^{\star}$s are shown in Table 1, for numbers in the main sweep.
$d_{\text{model}}$
$m$ $16$ $32$ $64$
$64$ 55 66 55
$128$ $11$ 2121 77
$256$ 66 $16$ $24$
$512$ 77 99 $14$
Table 1: Per-head key dimension $d_{k}$ from the main sweep.
These are found using the training step upper bounds shown in Table 2, where we increase the steps as the problem size and complexity increases.
$m$ $d_{\text{model}}$ Training step cutoff
64 16, 32, 64 20,000
128 16, 32, 64 20,000
256 32, 64 20,000
256 16 30,000
512 32, 64 20,000
512 16 80,000
1024 128 80,000
2048 256 80,000
4096 512 200,000
Table 2: Training step cutoffs by configuration. Default cutoff is 20,000 steps, with extended budgets for larger problem sizes.
Also, we provide additional examples of our findings from the main sweep of configurations in Fig. 6.
[…]
In Fig. 7, we plot the number of heads used in the optimal found configuration versus the compression $m/d_{\text{model}}$.
Figure 7: The number of heads needed grows approximately linearly with compression; the dashed line shows a least‑squares fit. See text for a description of the error bars. We do not tie the line to the origin, since heads are clipped at $h\geq 1$.
A.3 Sensitivity to context length.
[…]
Appendix B Further Details on Our Explicit Constructions
We here provide additional details on our explicit constructions from Section 4. We start with the easiest case - permutation graphs with no embedding. This result is subsummed by the second construction, so is only included as a warm up for the more general case.
B.1 Construction I: Permutation Graphs with One‑Hot Embeddings
Setup.
We start with the following two assumptions: (i) $G$ is a permutation on $m$ items, defined by a function $\pi:V\to V$, where the edges are $E=\{(v_{i},v_{\pi(i)})\mid v_{i}\in V\}$, and thus $m^{\prime}=m$. (ii) Node $v_{i}$ is represented by the one-hot vector $\mathbf{x}_{i}=\mathbf{e}_{i}\in\mathbb{R}^{m}$, setting the model dimension $d_{\text{model}}=m$.
With these assumptions, a single attention head ($h=1$) suffices, so the total key dimension is $D_{K}=d_{k}$. Our goal is to define $W_{K}$, $W_{Q}$, and a global threshold $\tau$ such that the score $S_{ij}=(\mathbf{x}_{i}W_{Q})\cdot(\mathbf{x}_{j}W_{K})$ exceeds $\tau$ iff $j=\pi(i)$. Our construction works for all vertices $V$ of the graph $G$, independent of the current context in our model; we formalize how this applies to specific contexts below.
Algorithmic Construction (one-hot case)
The core idea is to assign each node $v_{j}$ a random ”signature” via its key vector $\mathbf{k}_{j}$. The query vector $\mathbf{q}_{i}$ for $v_{i}$ is the signature of its target, $v_{\pi(i)}$. The dot product between vectors is maximized when the query signature matches the key signature.
Algorithm 2 Construction for Permutation Graphs with One-Hot Inputs
1:Input: Graph $G=(V,E)$ defined by permutation $\pi$.
2:Setup: Choose a probability $p\in(0,1/2)$, e.g., $p=1/4$, and dimension $d_{k}=C\log m$, for sufficiently large constant $C$.
3:
4:Construct Key Matrix: Draw $W_{K}\in\mathbb{R}^{m\times d_{k}}$ with i.i.d. entries $(W_{K})_{jl}\sim\text{Bernoulli}(p)$.
5:   For each node $v_{j}$, the key is $\mathbf{k}_{j}=\mathbf{e}_{j}W_{K}$.
6:
7:Construct Query Matrix: For each node $v_{i}$, set its query $\mathbf{q}_{i}=\mathbf{k}_{\pi(i)}$.
8:   This is equivalent to setting the $i$-th row of $W_{Q}$ to be the $\pi(i)$-th row of $W_{K}$.
9:
10:Set Threshold: $\tau=\frac{p+p^{2}}{2}\,d_{k}$.
Theorem B.1 (Single-head recognition under one-hot inputs).
Under the construction above, for $d_{k}=C\log m$ with $C$ sufficiently large (depending only on $p$), we have with probability at least $1-m^{-3}$ over the draw of $W_{K}$ that
$S_{i,\pi(i)}>\tau\quad\text{and}\quad S_{ij}<\tau\;\;\text{for all }i\in V,\,j\neq\pi(i).$
Hence a single attention head correctly identifies all edges of $G$.
Proof.
For $j=\pi(i)$,
$S_{i,\pi(i)}=\mathbf{k}_{\pi(i)}\cdot\mathbf{k}_{\pi(i)}\sim\text{Binomial}(d_{k},p)$
with mean $\mu_{1}=d_{k}p$. For $j\neq\pi(i)$,
$S_{ij}=\mathbf{k}_{\pi(i)}\cdot\mathbf{k}_{j}\sim\text{Binomial}(d_{k},p^{2})$
with mean $\mu_{2}=d_{k}p^{2}$. Take $\tau=\frac{\mu_{1}+\mu_{2}}{2}=\frac{p+p^{2}}{2}d_{k}$.
For the (lower) tail at the true edge, the Chernoff bound gives
$\Pr\!\big[S_{i,\pi(i)}\leq\tau\big]\;\leq\;\exp\!\Big(-\tfrac{\mu_{1}\delta_{1}^{2}}{2}\Big)\quad\text{where}\quad\delta_{1}=1-\tfrac{\tau}{\mu_{1}}=\tfrac{1-p}{2},$
so $\Pr[S_{i,\pi(i)}\leq\tau]\leq\exp\!\big(-\frac{d_{k}p(1-p)^{2}}{8}\big)$. For the (upper) tail at non-edges, the Chernoff bound yields
$\Pr\!\big[S_{ij}\geq\tau\big]\;\leq\;\exp\!\Big(-\tfrac{\mu_{2}\delta_{2}^{2}}{2+\delta_{2}}\Big)\quad\text{where}\quad\delta_{2}=\tfrac{\tau}{\mu_{2}}-1=\tfrac{1-p}{2p},$
hence $\Pr[S_{ij}\geq\tau]\leq\exp\!\big(-\frac{d_{k}\,p(1-p)^{2}}{2(1+3p)}\big)$. A union bound over the $m$ target pairs and the m(m1)m(m-1) non-target pairs gives a total failure probability
$m\,e^{-c_{1}d_{k}}+m^{2}e^{-c_{2}d_{k}}\quad\text{with}\quad c_{1}=\tfrac{p(1-p)^{2}}{8},\;c_{2}=\tfrac{p(1-p)^{2}}{2(1+3p)}.$
Choosing $d_{k}=C\log m$ with $C>\max\{3/c_{1},2/c_{2}\}$ makes this at most $m^{-3}$, establishing the simultaneous separation $S_{i,\pi(i)}>\tau>S_{ij}$ and correctness. ∎
Lemma 4.1 immediately now yields correctness for every context, independent of context length. Our lower bound from Section 7 for this case is $\Omega(\frac{m^{\prime}}{d_{\text{model}}}\log m)=\Omega(\log m)$. Our construction achieves an upper bound of $d_{k}=O(\log m)$, demonstrating that the bound is tight for this class of problems. Also note that the threshold proof is identical if softmax is used; see Appendix C.
B.2 Construction II: Permutations Under Compressive Embeddings
We next prove the correctness of Construction II, which follows from Theorem 4.2, restated here for convenience.
Theorem B.2 (Multi‑head recognition under Gaussian unit‑norm embeddings).
Assume the setup and construction of Algorithm 1 with $h=\frac{m}{d_{\mathrm{model}}}$ heads, per‑head dimension $d_{k}=C\log m$ for $C$ sufficiently large, and $\tau=\tfrac{1}{2}d_{k}$. If $d_{\mathrm{model}}\geq c_{0}\log m$ for a sufficiently large absolute constant $c_{0}$, then with probability at least $1-m^{-3}$ over the draw of $(X,W_{\mathrm{sig}})$,
$\forall i\in V\;\exists\,k\in[h]\text{ with }i\in V_{k}:\quad S^{(k)}_{i,\pi(i)}>\tau\quad\text{and}\quad S^{(k)}_{ij}<\tau\ \ \forall j\neq\pi(i).$
Consequently, max‑pooling over heads correctly recognizes all edges and $D_{K}=h\,d_{k}=\Theta\!\Big(\frac{m\log m}{d_{\mathrm{model}}}\Big).$
Proof.
Let $\mathbf{u}_{i}:=\mathbf{x}_{i}X^{\top}=\mathbf{e}_{i}(XX^{\top})$ and $\boldsymbol{\delta}_{i}:=\mathbf{u}_{i}-\mathbf{e}_{i}\in\mathbb{R}^{m}.$ Thus $\mathbf{u}_{i}$ is the ii‑th row of the Gram matrix $G:=XX^{\top}$; it satisfies $\mathbf{u}_{i}(i)=1$ and, for $j\neq i$, $\mathbf{u}_{i}(j)=\langle\mathbf{x}_{i},\mathbf{x}_{j}\rangle$. Fix a head $k$ and a source $i\in V_{k}$. As in Construction I, write
$\mathbf{q}^{(k)}_{i}=\mathbf{u}_{i}W^{\prime}_{Q,(k)},\qquad\mathbf{k}^{(k)}_{j}=\mathbf{u}_{j}W^{\prime}_{K,(k)}.$
Using $\mathbf{u}_{t}=\mathbf{e}_{t}+\boldsymbol{\delta}_{t}$ and the definitions of $W^{\prime}_{Q,(k)}$ and $W^{\prime}_{K,(k)}$, decompose, for any $j$,
$S^{(k)}_{ij}$ $=(\mathbf{u}_{i}W^{\prime}_{Q,(k)})\cdot(\mathbf{u}_{j}W^{\prime}_{K,(k)})$
$=\underbrace{\mathbf{w}_{\pi(i)}\cdot\mathbf{w}_{j}\cdot\mathbb{I}(j\in T_{k})}_{\text{Signal}}+\underbrace{\mathbf{w}_{\pi(i)}\cdot\textstyle\sum_{t\in T_{k}}\delta_{j,t}\mathbf{w}_{t}}_{N_{1}}$
$\quad+\underbrace{\Big(\sum_{s\in V_{k}}\delta_{i,s}\mathbf{w}_{\pi(s)}\Big)\cdot\mathbf{w}_{j}\cdot\mathbb{I}(j\in T_{k})}_{N_{2}}+\underbrace{\Big(\sum_{s\in V_{k}}\delta_{i,s}\mathbf{w}_{\pi(s)}\Big)\cdot\Big(\sum_{t\in T_{k}}\delta_{j,t}\mathbf{w}_{t}\Big)}_{N_{3}}.$
Signal here means the contribution that would remain under a perfect inverse (i.e., if $XX^{\top}=I$): $\mathbf{w}_{\pi(i)}\!\cdot\!\mathbf{w}_{j}\cdot\mathbb{I}(j\in T_{k})$. The Noise terms $N_{1},N_{2},N_{3}$ arise solely from the leakage vectors $\boldsymbol{\delta}_{i},\boldsymbol{\delta}_{j}$ due to approximate de-embedding. For $j\in T_{k}\setminus\{\pi(i)\}$ the cross-inner product $\mathbf{w}_{\pi(i)}\!\cdot\!\mathbf{w}_{j}$ is not counted as noise (it is intrinsic signature cross-correlation) and is bounded separately. To bound the Noise terms, we next quantify properties of the approximate inverse $XX^{\top}$ for unit‑norm Gaussian rows.
Lemma B.3 (Concentration of the approximate inverse).
Let $X$ be as above and $d_{\mathrm{model}}\geq c_{0}\log m$ for a sufficiently large constant $c_{0}$. With probability at least $1-m^{-4}$, simultaneously for all $i\in[m]$ and heads $k\in[h]$:
  1. 1.
    $\mathbf{u}_{i}(i)=1$ (deterministically).
  2. 2.
    (Leakage $L_{2}$‑mass) For $S\in\{V_{k}\setminus\{i\},\,T_{k}\}$,
    $\|\boldsymbol{\delta}_{i,S}\|_{2}^{2}=\sum_{s\in S}\langle\mathbf{x}_{i},\mathbf{x}_{s}\rangle^{2}\ \leq\ C_{2}$
    for an absolute constant $C_{2}$ (e.g., $C_{2}=2$).
  3. 3.
    (Cross‑correlations) For all $i,j$,
    $\Big|\sum_{a\in T_{k}}\delta_{i,\pi^{-1}(a)}\,\delta_{j,a}\Big|\ \leq\ C_{3}\,\sqrt{\frac{\log m}{d_{\mathrm{model}}}}$
    for an absolute constant $C_{3}$.
Proof.
For $j\neq i$, $\langle\mathbf{x}_{i},\mathbf{x}_{j}\rangle$ is mean‑zero sub‑Gaussian with parameter $\Theta(1/\sqrt{d_{\mathrm{model}}})$, and $\{\langle\mathbf{x}_{i},\mathbf{x}_{j}\rangle\}_{j\in S}$ are independent given $\mathbf{x}_{i}$. Then $(\langle\mathbf{x}_{i},\mathbf{x}_{j}\rangle^{2})_{j\in S}$ are i.i.d. sub‑exponential with $\psi_{1}$‑norm $\Theta(1/d_{\mathrm{model}})$ and mean $1/d_{\mathrm{model}}$. For $|S|=d_{\mathrm{model}}$, Bernstein’s inequality gives
$\Pr\!\Big[\sum_{s\in S}\!\langle\mathbf{x}_{i},\mathbf{x}_{s}\rangle^{2}>2\Big]\ \leq\ e^{-\Omega(d_{\mathrm{model}})}.$
A union bound over $i$ and the $2h$ choices of $S$ (recall $h=m/d_{\mathrm{model}}$) yields Item 2.
For Item 3, define independent mean‑zero sub‑exponential variables $Y_{a}:=\langle\mathbf{x}_{i},\mathbf{x}_{\pi^{-1}(a)}\rangle\cdot\langle\mathbf{x}_{j},\mathbf{x}_{a}\rangle$ for $a\in T_{k}$. Each has $\psi_{1}$‑norm $\Theta(1/d_{\mathrm{model}})$ and $\mathbb{E}Y_{a}=0$. Bernstein’s inequality implies $\Pr\big[\big|\sum_{a\in T_{k}}Y_{a}\big|\geq t\big]\leq 2\exp(-\Omega(\min\{d_{\mathrm{model}}t^{2},\ d_{\mathrm{model}}t\})).$ Taking $t=C_{3}\sqrt{(\log m)/d_{\mathrm{model}}}$ and union bounding over all $i,j,k$ proves Item 3 for $c_{0}$ large enough. Item 1 is immediate from unit‑norm rows. ∎
Signal.
If $j=\pi(i)$, then $\text{Signal}=\|\mathbf{w}_{\pi(i)}\|_{2}^{2}=d_{k}$ (exactly). If $j\in T_{k}$ and $j\neq\pi(i)$, then $\text{Signal}=\mathbf{w}_{\pi(i)}\cdot\mathbf{w}_{j}$ is a sum of $d_{k}$ i.i.d. Rademacher variables and thus sub‑Gaussian with mean 0 and variance $d_{k}$. By a union bound over all $(i,j,k)$, with probability at least $1-m^{-5}$,
$|\text{Signal}|\ \leq\ C_{\star}\sqrt{d_{k}\log m}\qquad\text{for all }(i,j\in T_{k}\setminus\{\pi(i)\},k),$
for an absolute constant $C_{\star}$.
Noise.
Condition on $X$ and apply Lemma B.3. For $N_{1}$,
$N_{1}=\sum_{r=1}^{d_{k}}\Big(\sum_{t\in T_{k}}\delta_{j,t}\,w_{t}[r]\Big)\,w_{\pi(i)}[r]$
is a sum of $d_{k}$ i.i.d. mean‑zero sub‑Gaussian variables with variance proxy $\|\boldsymbol{\delta}_{j,T_{k}}\|_{2}^{2}\leq C_{2}$. Hence, by Bernstein/Hoeffding and a union bound over $(i,j,k)$,
$|N_{1}|\ \leq\ C_{4}\sqrt{C_{2}\,d_{k}\,\log m}$
holds w.h.p. for an absolute constant $C_{4}$. The same bound holds for $N_{2}$ with $\|\boldsymbol{\delta}_{i,V_{k}}\|_{2}^{2}\leq C_{2}$.
For $N_{3}$, write for each column $r$,
$X_{r}:=\sum_{s\in V_{k}}\delta_{i,s}\,w_{\pi(s)}[r],\qquad Y_{r}:=\sum_{t\in T_{k}}\delta_{j,t}\,w_{t}[r].$
Then $N_{3}=\sum_{r=1}^{d_{k}}X_{r}Y_{r}$. Conditional on $X$, $\{(X_{r},Y_{r})\}_{r=1}^{d_{k}}$ are i.i.d.; each $X_{r}$ and $Y_{r}$ is mean‑zero sub‑Gaussian with parameters $\lesssim\|\boldsymbol{\delta}_{i,V_{k}}\|_{2}\leq\sqrt{C_{2}}$ and $\lesssim\|\boldsymbol{\delta}_{j,T_{k}}\|_{2}\leq\sqrt{C_{2}}$, respectively. Thus $X_{r}Y_{r}$ is mean $\langle\boldsymbol{\delta}_{i,\pi^{-1}(T_{k})},\boldsymbol{\delta}_{j,T_{k}}\rangle$ and sub‑exponential with $\psi_{1}$‑norm $\lesssim C_{2}$. Consequently,
$\mathbb{E}[N_{3}\mid X]\;=\;d_{k}\,\big\langle\boldsymbol{\delta}_{i,\pi^{-1}(T_{k})},\boldsymbol{\delta}_{j,T_{k}}\big\rangle,$
and, by Bernstein plus a union bound,
$\big|N_{3}-\mathbb{E}[N_{3}\mid X]\big|\ \leq\ C_{5}\,C_{2}\,\sqrt{d_{k}\log m}$
w.h.p. for an absolute constant $C_{5}$. Using Lemma B.3(3),
$\big|\mathbb{E}[N_{3}\mid X]\big|\ \leq\ d_{k}\,C_{3}\,\sqrt{\frac{\log m}{d_{\mathrm{model}}}}.$
Separation.
Choose constants $c_{0},C$ large enough so that
$C_{3}\,\sqrt{\tfrac{\log m}{d_{\mathrm{model}}}}\ \leq\ \tfrac{1}{16}\qquad\text{and}\qquad(C_{\star}+2C_{4}\sqrt{C_{2}}+C_{5}C_{2})\sqrt{\tfrac{\log m}{d_{k}}}\ \leq\ \tfrac{1}{16}.$
This is feasible since $d_{\mathrm{model}}\geq c_{0}\log m$ and $d_{k}=C\log m$.
Target edge $j=\pi(i)$. Using the bounds above (recall $\text{Signal}=d_{k}$ exactly),
$S^{(k)}_{i,\pi(i)}\;\geq\;d_{k}\ -\ \underbrace{(C_{4}\sqrt{C_{2}\,d_{k}\log m})}_{|N_{1}|}\ -\ \underbrace{(C_{4}\sqrt{C_{2}\,d_{k}\log m})}_{|N_{2}|}\ -\ \underbrace{\big(C_{5}C_{2}\sqrt{d_{k}\log m}+\tfrac{1}{16}d_{k}\big)}_{|N_{3}|}\;>\;\tfrac{3}{4}d_{k}\;>\;\tau.$
Non‑edge $j\neq\pi(i)$. If $j\notin T_{k}$ then $\text{Signal}=0$ and $N_{2}=0$, so
$|S^{(k)}_{ij}|\ \leq\ C_{4}\sqrt{C_{2}\,d_{k}\log m}+\big(C_{5}C_{2}\sqrt{d_{k}\log m}+\tfrac{1}{16}d_{k}\big)\ <\ \tfrac{1}{4}d_{k}\ <\ \tau.$
If $j\in T_{k}\setminus\{\pi(i)\}$, then $|\text{Signal}|\leq C_{\star}\sqrt{d_{k}\log m}$ and the same bounds for $N_{1},N_{2},N_{3}$ apply, giving $|S^{(k)}_{ij}|<\tfrac{1}{4}d_{k}<\tau$.
A union bound over all $(i,j,k)$ completes the proof. ∎
Thus, with Gaussian unit‑norm embeddings and Rademacher signatures, our construction recognizes the entire graph using a total key dimension
$D_{K}=h\cdot d_{k}=O\!\left(\frac{m\log m}{d_{\text{model}}}\right),$
This bound is asymptotically optimal, matching our lower bound within a constant factor. In the proof of Theorem 4.2, the non-edge bounds hold uniformly over all $(i,j,k)$ (we union bound over $(i,j,k)$), so for any $j\neq\pi(i)$ we have $S_{ij}^{(k)}<\tau$ for all heads $k$ and hence $S_{ij}^{\max}<\tau$, while the target edge satisfies $S_{i,\pi(i)}^{\max}>\tau$. Lemma 4.1 then yields correctness on $E|_{\mathcal{C}}$ for every context $\mathcal{C}$.
B.3 Construction III: More General Embeddings
The analysis of Construction II (Gaussian unit–norm) ultimately used only two facts about the Gram matrix $XX^{\top}$: (i) diagonals concentrate around a common scale, and (ii) for any small subset of indices the off–diagonal leakage has bounded $\ell_{2}$ mass, with a mild control on a corresponding cross–leakage term. We package these into a reusable, block–level notion that subsumes the usual pairwise incoherence and is tight enough to cover sparse/binary compressive embeddings.
Definition B.4 (Restricted self–incoherence at block size $B$).
Fix parameters $\mu>0$, $\varepsilon_{d}\in[0,1)$, block size $B\in\mathbb{N}$, and leakage levels $\rho,\gamma\geq 0$. An embedding matrix $X\in\mathbb{R}^{m\times d_{\text{model}}}$ with rows $\{\mathbf{x}_{i}\}_{i=1}^{m}$ is $(\mu,\varepsilon_{d},B;\rho,\gamma)$–restricted self–incoherent if, writing
$X_{\mathrm{inv}}:=\frac{1}{\mu}X^{\top},\qquad\mathbf{u}_{i}:=\mathbf{x}_{i}X_{\mathrm{inv}}=\frac{1}{\mu}\mathbf{e}_{i}(XX^{\top}),\qquad\boldsymbol{\delta}_{i}:=\mathbf{u}_{i}-\mathbf{e}_{i},$
the following hold simultaneously:
  1. 1.
    Diagonal stability: $\mathbf{u}_{i}(i)\in[1-\varepsilon_{d},\,1+\varepsilon_{d}]$ for all $i$.
  2. 2.
    Restricted leakage mass: for every $i$ and every $S\subseteq[m]\setminus\{i\}$ with $|S|\leq B$,
    $\|\boldsymbol{\delta}_{i,S}\|_{2}^{2}=\sum_{s\in S}\delta_{i}(s)^{2}\ \leq\ \rho.$
  3. 3.
    Restricted cross–leakage: for every $i,j$ and $S\subseteq[m]$ with $|S|\leq B$,
    $\Big|\sum_{a\in S}\delta_{i}(a)\,\delta_{j}(a)\Big|\ \leq\ \gamma.$
Algorithm 3 Construction for Generalized Embeddings
1:Input: Embedding matrix $X\in\mathbb{R}^{m\times d_{\text{model}}}$; permutation graph $G=(V,E)$ with $\pi:V\to V$.
2:Parameters: Signature sparsity $p\in(0,1/20]$; per–head width $d_{k}=C\log m$ for a sufficiently large absolute constant $C$.
3:Random signatures: Draw $W_{\mathrm{sig}}\in\{0,1\}^{m\times d_{k}}$ with i.i.d. Bernoulli(p)(p) entries; let $\mathbf{w}_{j}$ denote its jj‑th row.
4:Set Threshold: $\tau:=\tfrac{p+p^{2}}{2}\,d_{k}$.
5:Choose block size and partition: Pick a block size $B$ (specified per embedding family below). Let $h:=\lceil m/B\rceil$ and partition $V$ into blocks $V_{1},\dots,V_{h}$ with $|V_{k}|\leq B$. For each head $k$, define its target set
$T_{k}\;:=\;\{\pi(s):\,s\in V_{k}\}.$
6:Define one‑hot–space templates (for each head $k$):
$W^{\prime}_{Q,(k)}(i,:)\;=\;\begin{cases}\mathbf{w}_{\pi(i)}&i\in V_{k}\\[2.0pt] 0&\text{else}\end{cases},\qquad W^{\prime}_{K,(k)}(j,:)\;=\;\begin{cases}\mathbf{w}_{j}&j\in T_{k}\\[2.0pt] 0&\text{else}\end{cases}.$
7:Realize parameters via approximate inverse:
$W^{(k)}_{Q}\;=\;X_{\mathrm{inv}}\,W^{\prime}_{Q,(k)},\qquad W^{(k)}_{K}\;=\;X_{\mathrm{inv}}\,W^{\prime}_{K,(k)}.$
Theorem B.5 (Recognition under restricted self–incoherence).
Let $X$ be $(\mu,\varepsilon_{d},B;\rho,\gamma)$–restricted self–incoherent for some $B$. Fix any $p\leq 1/20$, take $d_{k}=C\log m$ with $C$ a sufficiently large absolute constant, and set $\tau=\tfrac{p+p^{2}}{2}d_{k}$. There exist absolute numerical constants $(c_{1},c_{2},c_{3})$ such that if
$\varepsilon_{d}\leq c_{1},\qquad\rho\ \leq\ \frac{c_{2}}{\log m},\qquad\gamma\ \leq\ \frac{c_{3}}{\log m},$
then with probability at least $1-m^{-3}$ over $W_{\mathrm{sig}}$ (and the draw of $X$ if random),
$\forall i\in V\;\exists\,k\in[h]\text{ with }i\in V_{k}:\quad S^{(k)}_{i,\pi(i)}>\tau\quad\text{and}\quad S^{(k)}_{ij}<\tau\ \ \forall j\neq\pi(i).$
Consequently, max–pooling over heads recovers all edges and the total key budget satisfies
$D_{K}=h\,d_{k}=\Theta\!\Big(\frac{m\log m}{B}\Big).$
Proof sketch.
As in Construction II, the score decomposes into a signal term plus three noise terms: $S^{(k)}_{ij}=\mathbf{w}_{\pi(i)}\!\cdot\!\mathbf{w}_{j}\cdot\mathbb{I}(j\in T_{k})+N_{1}+N_{2}+N_{3},$ with $N_{1},N_{2},N_{3}$ arising from $\boldsymbol{\delta}_{i},\boldsymbol{\delta}_{j}$. Write $\mathbf{u}_{t}=\mathbf{e}_{t}+\boldsymbol{\delta}_{t}$ and expand $S^{(k)}_{ij}$ as in Construction II. Conditioned on $X$, each column of $W_{\mathrm{sig}}$ contributes an independent copy of the signal/noise decomposition. Using Chernoff for the Bernoulli signal coordinates gives, uniformly over all $(i,j,k)$, the standard separation $\mu_{1}-\mu_{2}=(p-p^{2})d_{k}$ between $j=\pi(i)$ and $j\in T_{k}\setminus\{\pi(i)\}$ up to $O(\sqrt{d_{k}\log m})$ fluctuations.
For $N_{1}$ and $N_{2}$, restricted leakage mass yields $\mathrm{Var}(N_{1}),\mathrm{Var}(N_{2})\ \lesssim\ d_{k}\,\rho$ and hence $|N_{1}|,|N_{2}|\ \lesssim\ \sqrt{d_{k}\,\rho\,\log m}$ uniformly with probability $1-m^{-5}$. For $N_{3}$, the centered part concentrates at scale $\lesssim\sqrt{d_{k}\,\log m}\cdot(\rho)^{1/2}$, while the mean shift equals $d_{k}\langle\boldsymbol{\delta}_{i,\pi^{-1}(T_{k})},\boldsymbol{\delta}_{j,T_{k}}\rangle$ and is controlled by $\gamma$. Choosing $C$ large and $(c_{1},c_{2},c_{3})$ small makes the total noise $\tfrac{1}{4}(p-p^{2})d_{k}$ uniformly, while the target signal sits $\tfrac{3}{4}(p-p^{2})d_{k}$ above $\mu_{2}$, giving the stated threshold separation. ∎
How to pick $B$.
The theorem asks only that $\rho,\gamma\lesssim 1/\log m$ at the chosen block size $B$. Different embedding families admit different $(\rho,\gamma)$–vs–$B$ trade–offs; plugging the corresponding $B$ into $D_{K}=\Theta((m/B)\log m)$ yields the budget.
Corollaries for common embedding models
Corollary B.6 (Gaussian unit–norm (GUN)).
Let each row $\mathbf{x}_{i}$ be drawn i.i.d. as $\tilde{\mathbf{x}}_{i}\sim\mathcal{N}(0,I/d_{\text{model}})$ and then $\ell_{2}$–normalized. Then w.h.p.
$\varepsilon_{d}=0,\qquad\rho\ \lesssim\ \frac{B}{d_{\text{model}}},\qquad\gamma\ \lesssim\ \sqrt{\frac{B}{d_{\text{model}}}},$
and Theorem B.5 holds for any $B\leq c\,d_{\text{model}}/\log m$. Choosing $B=\Theta(d_{\text{model}})$ yields
$h=\Theta\!\Big(\frac{m}{d_{\text{model}}}\Big),\qquad D_{K}=\Theta\!\Big(\frac{m\log m}{d_{\text{model}}}\Big),$
in agreement with Theorem 4.2 up to constants (the specialized proof in Construction II attains this with the sharp choice $B=d_{\text{model}}$).
Corollary B.7 (Random binary compressive embeddings (RBCE)).
[…]
$\mathbf{u}_{i}(i)\in[1-\varepsilon_{d},1+\varepsilon_{d}]\ \text{ with }\ \varepsilon_{d}\lesssim\frac{1}{\sqrt{\mu}},\qquad\rho\ \lesssim\ \frac{B}{d_{\text{model}}},\qquad\gamma\ \lesssim\ B\,p_{B}^{2}.$
Consequently, taking
$B\ =\ \Theta\!\Big(\frac{d_{\text{model}}}{\log m}\Big)\quad\Longrightarrow\quad\rho\ \lesssim\ \frac{1}{\log m},\ \ \gamma\ \lesssim\ \frac{\log m}{d_{\text{model}}},$
and Theorem B.5 applies. The number of heads and total key budget become
$h=\Theta\!\Big(\frac{m\log m}{d_{\text{model}}}\Big),\qquad D_{K}=h\,d_{k}=\Theta\!\Big(\frac{m\log^{2}m}{d_{\text{model}}}\Big).$
Proof idea for Corollary B.7.
[…]
$\sum_{s\in S}\!\langle\mathbf{x}_{i},\mathbf{x}_{s}\rangle^{2}\ \leq\ \sum_{s\in S}\!\langle\mathbf{x}_{i},\mathbf{x}_{s}\rangle\quad\text{and}\quad\mathbb{E}\big[\langle\mathbf{x}_{i},\mathbf{x}_{s}\rangle\big]=d_{\text{model}}p_{B}^{2},$
so $\mathbb{E}\|\boldsymbol{\delta}_{i,S}\|_{2}^{2}=\frac{1}{\mu^{2}}\sum_{s\in S}\mathbb{E}\langle\mathbf{x}_{i},\mathbf{x}_{s}\rangle^{2}\lesssim B/d_{\text{model}}$, and a Bernstein + union bound yields $\rho\lesssim B/d_{\text{model}}$. Similarly, $\mathbb{E}\sum_{a\in S}\delta_{i}(a)\delta_{j}(a)=\frac{|S|}{\mu^{2}}\,\mathbb{E}\langle\mathbf{x}_{i},\mathbf{x}_{a}\rangle\,\mathbb{E}\langle\mathbf{x}_{j},\mathbf{x}_{a}\rangle\lesssim Bp_{B}^{2},$ and concentration gives $\gamma\lesssim Bp_{B}^{2}$ uniformly. ∎
Signature family.
We stated the construction with Bernoulli(p)(p) signatures because the thresholding analysis naturally separates $j=\pi(i)$ from $j\in T_{k}\setminus\{\pi(i)\}$ at means $pd_{k}$ vs. $p^{2}d_{k}$. One can equivalently use Rademacher $\{\pm 1\}$ signatures with threshold $\tau=\tfrac{1}{2}d_{k}$; all bounds above translate verbatim with the same $B$ and $d_{k}=\Theta(\log m)$.
Takeaways.
Definition B.4 abstracts the only geometric inputs needed by the attention construction. Plugging in model–specific $(\rho,\gamma)$–vs–$B$ trade–offs yields the head count $h=\Theta(m/B)$ and total key budget $D_{K}=\Theta((m/B)\log m)$. For Gaussian unit–norm embeddings one recovers the $D_{K}=\Theta(m\log m/d_{\text{model}})$ guarantee; for sparse random binary compressive embeddings one obtains $D_{K}=\Theta(m\log^{2}m/d_{\text{model}})$.
B.4 Construction IV: General Graphs
We now extend the permutation constructions to general directed graphs $G=(V,E)$ with $|V|=m$ vertices and $|E|=m^{\prime}$ edges. In this case, our information theoretic lower bound on total key dimension is $D_{K}=\Omega\!\left(\frac{m^{\prime}}{d_{\text{model}}}\,\log(m^{2}/m^{\prime})\right).$ We here provide a general upper bound for any graph, and show that for graphs that have a mild skew condition (the maximum degree is not too much larger than the average degree), it asymptotically matches this lower bound for all but the densest graphs (which match within a log factor). As before we use max aggregation over heads with a global scalar threshold $\tau$, and we work under the Gaussian unit‑norm embedding model from Construction II: the row vectors of $X\in\mathbb{R}^{m\times d_{\text{model}}}$ are i.i.d. isotropic Gaussian followed by $L_{2}$-normalization. All probabilities are over the draw of $X$ and of the (head-shared) random signature matrix.
Packing edges into matchings.
The analysis in Theorem 4.2 operates on blocks in which each source has exactly one outgoing edge and targets are distinct within the block. Equivalently, each head should see a matching (a partial permutation) between a set of sources and a set of targets.
We will use a simple decompositions of the edge set into matchings of size $d_{\text{model}}$ which will be our block size. Write $d_{\mathrm{out}}(i)$ and $d_{\mathrm{in}}(i)$ for the out-/in-degree of $v_{i}$. Let $\Delta_{\mathrm{out}}:=\max_{i}d_{\mathrm{out}}(i)$ and $\Delta_{\mathrm{in}}:=\max_{i}d_{\mathrm{in}}(i)$ denote the maximum out- and in-degrees, and write $\Delta:=\max\{\Delta_{\mathrm{out}},\Delta_{\mathrm{in}}\}$.
Lemma B.8 (Coloring-and-batching decomposition).
Let $G=(V,E)$ be any directed graph on $m$ vertices and $m^{\prime}$ edges, and let $H:=\left\lceil\frac{m^{\prime}}{d_{\text{model}}}\right\rceil+\Delta$. Then there exists a partition of $E$ into $H$ disjoint sets $M_{1},\dots,M_{H}$ such that for every $k$: (i) $M_{k}$ is a matching (no two edges in $M_{k}$ share a source or a target); (ii) $|M_{k}|\leq d_{\text{model}}$.
Proof.
Identify $G$ with its bipartite incidence graph $\mathcal{B}=(V_{L}\!\cup\!V_{R},E)$ where each directed edge $(i,j)$ becomes an undirected edge between $i\in V_{L}$ and $j\in V_{R}$. Then $\Delta(\mathcal{B})=\Delta$. By Kőnig’s line‑coloring theorem, $E=F_{1}\cup\cdots\cup F_{\Delta}$ with each $F_{c}$ a matching. Split each $F_{c}$ into blocks of size at most $d_{\text{model}}$; since $\sum_{c=1}^{\Delta}\lceil|F_{c}|/d_{\text{model}}\rceil\leq\left\lceil\frac{\sum_{c}|F_{c}|}{d_{\text{model}}}\right\rceil+\Delta=\left\lceil\frac{m^{\prime}}{d_{\text{model}}}\right\rceil+\Delta=H$, we obtain $H$ matchings $M_{k}$ each of size at most $d_{\text{model}}$. ∎
Thus, after packing via Lemma B.8 head $k$ will operate on the matching $M_{k}$. Let $V_{k}\subseteq V$ and $T_{k}\subseteq V$ denote the sources and targets incident to $M_{k}$ and write $\pi_{k}:V_{k}\to T_{k}$ for the bijection defined by $M_{k}$.
Construction.
We reuse the compressive permutation machinery head‑by‑head.
Algorithm 4 Construction for General Graphs
1:Input: Directed graph $G=(V,E)$ with $|V|=m$, $|E|=m^{\prime}$; embedding matrix $X\in\mathbb{R}^{m\times d_{\text{model}}}$ with Gaussian unit‑norm rows.
2:Parameters: Number of heads $h=H=\left\lceil\frac{m^{\prime}}{d_{\text{model}}}\right\rceil+\Delta$; per‑head key/query dimension $d_{k}=C\log m$ for a sufficiently large absolute constant $C$. Each head uses block size $d_{\text{model}}$.
3:Pack edges into matchings: Decompose $E$ into disjoint matchings $M_{1},\dots,M_{H}$ with $|M_{k}|\leq d_{\text{model}}$ using Lemma B.8. For each $k$, let $V_{k}$ and $T_{k}$ be the sources and targets incident to $M_{k}$ and write $\pi_{k}:V_{k}\to T_{k}$ for the associated bijection.
4:Random signatures: Draw a shared Rademacher matrix $W_{\mathrm{sig}}\in\{\pm 1\}^{m\times d_{k}}$ with i.i.d. entries; let $\mathbf{w}_{j}$ denote its jj‑th row.
5:Per‑head “ideal” matrices:
$\big(W^{\prime}_{Q,(k)}\big)_{i,\cdot}:=\begin{cases}\mathbf{w}_{\pi_{k}(i)}&i\in V_{k}\\ \mathbf{0}&\text{otherwise}\end{cases},\qquad\big(W^{\prime}_{K,(k)}\big)_{j,\cdot}:=\begin{cases}\mathbf{w}_{j}&j\in T_{k}\\ \mathbf{0}&\text{otherwise}\end{cases},$
where $\mathbf{w}_{j}$ is the jj‑th row of $W_{\mathrm{sig}}$.
6:Final projections (approximate de‑embedding): As in Construction II, use the approximate inverse $X^{\top}$:
$W_{Q}^{(k)}\;=\;X^{\top}W^{\prime}_{Q,(k)},\qquad W_{K}^{(k)}\;=\;X^{\top}W^{\prime}_{K,(k)}.$
7:Set Threshold: $\tau=\frac{1}{2}d_{k}$.
Theorem B.9 (General graphs).
Assume $d_{\text{model}}\geq c_{0}\log m$ for a sufficiently large constant $c_{0}$. With the construction above (using $h=\left\lceil\frac{m^{\prime}}{d_{\text{model}}}\right\rceil+\Delta$ heads and $d_{k}=C\log m$), there is a universal $C$ such that, with probability at least $1-m^{-3}$ over the draw of $(X,W_{\mathrm{sig}})$, simultaneously for all ordered pairs $(i,j)$,
$S_{ij}^{\max}\;=\;\max_{1\leq k\leq H}S_{ij}^{(k)}\begin{cases}>\ \tau&\text{if }(i,j)\in E,\\ <\ \tau&\text{if }(i,j)\notin E.\end{cases}$
Consequently, $D_{K}=O\!\bigl(\frac{m^{\prime}\log m}{d_{\text{model}}}+\Delta\log m\bigr).$
Proof sketch.
By Lemma B.8, each head $k$ sees a matching $M_{k}$ of size at most $d_{\text{model}}$, with a bijection $\pi_{k}:V_{k}\to T_{k}$. Within head $k$, the score decomposition and concentration bounds are exactly those of Theorem 4.2: for $(i,j)=(i,\pi_{k}(i))$ the Signal term equals $d_{k}$ and the three Noise terms ($N_{1},N_{2},N_{3}$) are controlled using Lemma B.3, since all leakage sets ($V_{k}\setminus\{i\}$ and $T_{k}$) have size $\leq d_{\text{model}}$. For $(i,j)\neq(i,\pi_{k}(i))$, Signal is a sum of i.i.d. Rademachers with variance $d_{k}$, while the same leakage bounds control $N_{1},N_{2},N_{3}$. Choosing $C$ and $c_{0}$ as in Theorem 4.2 yields, within each head, $S_{i,\pi_{k}(i)}^{(k)}>\tau$ and $|S_{ij}^{(k)}|<\tau$ for all $j\neq\pi_{k}(i)$ simultaneously with probability $1-m^{-4}$.
A union bound over all heads and all pairs in those heads costs only a $\log$ factor absorbed by $d_{k}=C\log m$: using $|M_{k}|\leq d_{\text{model}}$ and $\sum_{k}|M_{k}|=m^{\prime}$, we have $\sum_{k}|M_{k}|^{2}\leq d_{\text{model}}\sum_{k}|M_{k}|=d_{\text{model}}m^{\prime}=O(m^{\prime}d_{\text{model}})$ events in total. Finally, max pooling across heads preserves separation (non‑edges are below $\tau$ in every head, and each true edge belongs to exactly one $M_{k}$), and Lemma 4.1 yields context‑robustness for arbitrary subsets $\mathcal{C}\subseteq V$. ∎
Degree skew and tightness.
Let $d_{\mathrm{avg}}=\frac{m^{\prime}}{m}$. Define the skew factor to be $\Delta/d_{\mathrm{avg}}$ and consider the condition
$\frac{\Delta}{d_{\mathrm{avg}}}\;\leq\;\frac{m}{d_{\text{model}}}.$ (10)
In other words, the ratio of the maximum degree to the average degree is no larger than the compression of the embedding, or equivalently, $\Delta\leq\frac{m^{\prime}}{d_{\text{model}}}$. This condition automatically holds for all dd‑regular graphs (since $\Delta=d_{\mathrm{avg}}$).
Corollary B.10 (Bounded Skew).
Assume $d_{\text{model}}\!\geq\!c_{0}\log m$ and (10). Then the construction with $h_{0}=\lceil m^{\prime}/d_{\text{model}}\rceil$ heads and $d_{k}=C\log m$ achieves the same separation guarantee as Theorem B.9, and $D_{K}=\Theta\!\bigl(\frac{m^{\prime}\log m^{\prime}}{d_{\text{model}}}\bigr).$
This is immediate from from Theorem B.9 and asymptotically matches the lower bound from Section 7 for this class of graphs, provided $m^{\prime}=O(m^{2-\epsilon})$ for some positive constant $\epsilon$.
Appendix C Additional Justification for the Model (Section 3)
Computational footprint.
While it is natural to consider the number of heads ($h$) and the per-head key/query dimension ($d_{k}$) as two separate resources, we argue that the most relevant complexity measure is their product. In practice, the computation for multiple heads is not performed as $h$ distinct operations but as a single, larger batched operation. Let $\mathbf{X}\in\mathbb{R}^{\ell\times d_{\text{model}}}$ stack the context embeddings. With $W_{Q}^{\text{cat}},W_{K}^{\text{cat}}\in\mathbb{R}^{d_{\text{model}}\times(hd_{k})}$ formed by concatenating head weights, queries/keys are $\mathbf{Q}_{\text{total}}=\mathbf{X}W_{Q}^{\text{cat}}$ and $\mathbf{K}_{\text{total}}=\mathbf{X}W_{K}^{\text{cat}}$. Both flops and parameter/memory cost scale as $O(\ell\,d_{\text{model}}\,hd_{k})$ and $O(d_{\text{model}}\,hd_{k})$, respectively, motivating $D_{K}=hd_{k}$ as the budget.
While sub-cubic matrix multiplication algorithms could theoretically make one large head asymptotically faster than several smaller ones, this effect is absent in practice. The true bottleneck for these operations on modern hardware is memory bandwidth—the rate at which the matrices can be fetched from memory. The total data moved is proportional to the size of the weight matrices, which is in turn proportional to $d_{\text{model}}\cdot(h\cdot d_{k})$. Because deep learning libraries are highly optimized to perform these batched multiplications, the performance typically tracks the total size of the key and query matrices, regardless of the number of heads.
Analyzing the QK channel in isolation.
Since RGR asks where a source should connect, the OV pathway only reweights or propagates information after the routing decision has been made. Thus, a correct edge can only dominate if the QK gate already concentrates sufficient mass on the true target. Increasing value dimension $D_{V}$ amplifies whatever QK selects; it does not fix mis‑routing. [^5] Note that our model is scoped to a single self-attention layer; multi-layer iterative routing is outside our abstraction. Furthermore, our construction aligns with recent mechanistic analyses that separate each attention head into an OV circuit (what is read/written) and a QK circuit (where to attend); the QK circuit determines the attention pattern and thus the directed edges in RGR [33, 19].
Aggregating across heads.
This ‘OR‑of‑heads’ max is an analysis device for the edge test; it does not assert cross‑head QK interaction in standard MHA implementations (where heads are combined in the value pathway). However, aggregating multi-head QK scores by a max implements the intended RGR semantics: each head $k$ specializes to a relational template, and an (untyped) edge should exist if any template fires, i.e., $S^{\max}_{pq}=\max_{k}S^{(k)}_{pq}>\tau$ (an OR-of-relations). If a smooth alternative is desired, replacing $\max$ with log-sum-exp preserves the binary decision up to a global threshold shift, since for any scores $a_{1},\dots,a_{h}$,
$\max_{k}a_{k}\;\leq\;\operatorname{LSE}(a)=\log\!\sum_{k=1}^{h}e^{a_{k}}\;\leq\;\max_{k}a_{k}+\log h,$
so one can retune $\tau$ by at most $\log h$ without changing the classifier [6].
From a detection viewpoint, $\max$ is the standard OR-of-detectors aggregator (as in max-pooling): it passes through the strongest localized evidence while suppressing clutter, whereas averaging or summing allows many weak, unrelated heads to “conspire” to cross the threshold—undesirable for a yes/no edge test. This aligns with empirical observations that only a small number of specialized heads dominate while many can be pruned with little effect; a max aggregator naturally yields a principled winner-take-all over these specialists without paying for redundant heads.
Thresholding vs. Softmax.
Our model makes binary edge decisions from the raw QK scores $S_{ij}$. As a result, our thresholding decision paradigm can be replaced with the usual $1/\sqrt{d_{k}}$ scaling and softmax without changing the asymptotic budgets used by our constructions for all but very dense graphs. Throughout this paper, we keep the thresholding rule for simplicity and ease of exposition, but we describe how to incorporate this change into our results.
Scaling. If the decision is made directly on raw scores, the $1/\sqrt{d_{k}}$ factor can be absorbed into the threshold on $S_{ij}$. If the decision is made after softmax, we absorb $1/\sqrt{d_{k}}$ into the learned projections $W_{Q},W_{K}$ so that the effective logits are unchanged.
Softmax. Let $a_{ij}=\exp(S_{ij})/\sum_{t\in\mathcal{C}}\exp(S_{it})$ be the row-wise softmax over a context $\mathcal{C}$ (we assume distinct members of $V$ in any context, so $|\mathcal{C}|=\ell\leq m$). We use a single global (per graph) threshold $\hat{\tau}$ on the attention weight:
$\text{decide $(i,j)$ is an edge}\quad\Longleftrightarrow\quad a_{ij}\ \geq\ \hat{\tau}\quad\Longleftrightarrow\quad S_{ij}-\operatorname{lse}_{t\in\mathcal{C}}S_{it}\ \geq\ \log\hat{\tau}$.
Let $\Delta$ be the maximum out-degree of the graph, and let
$\gamma\ :=\ \min_{i}\Big(\min_{j\in N_{i}}S_{ij}\ -\ \max_{j\notin N_{i}}S_{ij}\Big)$
be the uniform score gap. A standard bound [6] yields, for any positive $j\in N_{i}$ and any $\ell\leq m$,
$a_{ij}\ \geq\ \frac{1}{\,\Delta+(\ell-\Delta)e^{-\gamma}\,}.$ (11)
Therefore a single global threshold $\hat{\tau}$ works for all contexts $\ell\leq m$ provided
$\gamma\ \geq\ \log\!\Big(\frac{m-\Delta}{\,1/\hat{\tau}-\Delta\,}\Big)\qquad\text{and necessarily}\qquad\hat{\tau}\ \leq\ \frac{1}{\Delta}.$ (12)
The condition $\hat{\tau}\leq 1/\Delta$ is unavoidable because the $k$ positive weights in a row sum to at most $1$. If $\Delta\leq m^{1-\varepsilon}$ for some constant $\varepsilon\in(0,1]$ and we take $\hat{\tau}=c/\Delta$ with any fixed $c\in(0,1)$, then
$\gamma\ \geq\ \log\!\Big(\frac{m-\Delta}{\,\Delta(1/c-1)\,}\Big)\ =\ \log\!\Big(\tfrac{m}{\Delta}-1\Big)\;+\;\log\!\Big(\tfrac{c}{1-c}\Big)\ \geq\ \varepsilon\log m\;+\;O(1).$
In all of our signature-based constructions, the gap satisfies $\gamma=\Theta(d_{k})$ with high probability; hence choosing $d_{k}=C\log m$ with $C$ large enough to meet (12) yields a valid single global threshold $\hat{\tau}$ for all contexts $\ell\leq m$.
We also note that sparse alternatives to softmax (sparsemax and $\alpha$-entmax) implement threshold-like attention rules that align directly with our edge test [43, 48, 11].
Appendix D Related Work
We survey work most relevant to our capacity-centric view of self-attention and position our Relational Graph Recognition (RGR) results in that landscape. The central distinction we draw is between what attention can compute in principle (expressivity), how architectural resources govern this power (capacity), and which parts of the Transformer carry the binding/addressing load (keys/queries vs. other channels).
Memorization Capacity and Parameter–Dependent Bounds
A growing body of work quantifies how many input–label associations Transformers—and, more narrowly, the attention mechanism—can memorize. As described in Section 2, the bounds from the memorization setting do not directly imply bounds on RGR, nor the other way around. For the attention module itself, [42] prove that a single MHA layer with $h$ heads can memorize $\Omega\!\big(h\,\min\{\ell,d_{k}\}\big)$ examples under a linear-independence assumption on the inputs, highlighting linear scaling in $h$ and the role of the per-head key/query width $d_{k}$. Complementary analyses bound attention’s memory depth and clarify depth–capacity trade-offs [41]. Moving to full Transformers, constructive results show that (under token-wise $(r,\delta)$-separated inputs) a stack of $2\ell$ self-attention layers suffices to memorize $N$ sequences with $\tilde{O}\!\big(\ell+\sqrt{\ell N}\big)$ parameters [35]; even a single-layer, single-head Transformer has nontrivial capacity under the same separatedness assumption, whereas replacing softmax by hardmax breaks memorization [31].
Beyond construction-style bounds, [41] give general upper and lower bounds for next-token prediction that scale as $\Theta(\omega N)$ in the presence of positional encodings and a vocabulary of size $\omega$, and [8] show that a single-layer Transformer can memorize when sequences are sufficiently zero-padded (though not in a parameter-optimal way). Classical results for ReLU networks connect parameter counts to memorization thresholds and VC-style capacity [58, 59]. More closely related to our focus on resource efficiency, [32] establish nearly matching upper/lower bounds on the minimal parameter count needed for memorization in Transformers: $\tilde{O}(\sqrt{N})$ parameters are sufficient (and necessary up to logs) for next-token prediction, and $\tilde{O}(\sqrt{\ell N})$ for sequence-to-sequence, under token-wise separatedness; they further suggest that self-attention effectively identifies sequences while the feed-forward network can become the bottleneck when associating labels.
Superposition, Constructive Designs, and Depth Separation
A concurrent line of work analyzes how networks compute many features in superposition, with lower and upper bounds for narrow MLPs and constructive designs for multi-feature computation [2, 1, 26]. The capacity limits shown there (stated as upper and lower bounds on neurons to compute a number of Boolean functions in parallel) are complementary to ours in terms of architectures: their focus is on MLPs while ours is on attention.
Foundational depth-separation and minimal-width universality results motivate the proof template we adopt—information-theoretic lower bounds matched by explicit constructions [55, 25, 34, 12]. In attention, constructive correspondences also explain how multi-head architectures partition pattern spaces; e.g., with relative positions, $s^{2}$ heads can realize any $s\times s$ convolution [10]. Our constructions similarly partition relational signal across heads to mitigate interference when $d_{\text{model}}\!\ll\!m$, explaining the empirical advantage of many small heads and clarifying when too-small $d_{k}$ triggers low-rank failure [5].
Dimension-, Rank-, and Resource–Driven Expressivity
A growing body of theory isolates how dimensional resources govern attention’s representational power. Universality guarantees establish that sufficiently resourced Transformers can approximate sequence-to-sequence functions [70], while more refined results show task-dependent strengths and weaknesses [52]. Focusing on the attention map, [39] prove that with fixed error and sparsity, self-attention can approximate dynamic sparse right-stochastic matrices using only $O(\log\ell)$ hidden dimensions (for context length $\ell$), echoing the role of near-orthogonality we exploit in our constructions. Conversely, [5] identify a per-head low-rank bottleneck: when $d_{k}<\ell$, a head cannot realize arbitrary $\ell\times\ell$ stochastic attention matrices. This clarifies a trade-off inside the total key/query budget $D_{K}=h\,d_{k}$: pushing $D_{K}$ into many tiny heads can induce head-wise rank limits.
Beyond these, several works develop structural and inductive-bias characterizations of self-attention. [15] show that pure attention without mixing loses rank doubly-exponentially with depth, explaining failure modes in deep attention stacks and underscoring the role of residual mixing. [17] analyze variable creation and sparsity patterns induced by softmax, while [51] use convex duality to give optimization- and geometry-based interpretations of ViT attention. For sample complexity and approximation, [37] study learning and generalization of shallow ViTs; rates and approximation guarantees have been developed for Transformer encoders and sequence models [23, 54, 30]. Recent generalization bounds that are (largely) sequence-length independent sharpen this picture [56]. Finally, theory has also pinpointed sparsity-oriented inductive biases: Transformers provably learn sparse token selection that FCNs cannot [68], and exhibit a simplicity bias for sparse Boolean functions [4].
Empirical observations likewise single out the key/query channel as an operative budget. Our results formalize this perspective for a concrete family (RGR), deriving matching lower and constructive upper bounds in terms of $D_{K}$ and the number of relations.
Formal-Language Limits, Compositionality, and Universality
Formal-language analyses delimit what fixed-size attention can recognize. Beyond general universality [70], there are sharp impossibility results for periodic and hierarchical languages [24, 3, 69]. Recent work uses communication-complexity arguments to show single-layer self-attention struggles with function composition at fixed embedding/heads, e.g., “grandparent-of” requires resources that scale with domain size [47]. Complementing these, [40] identify additional structural constraints on what Transformers can compute under realistic resource regimes. We view these results as orthogonal to RGR: they characterize classes of computations, whereas we fix a relational family and ask how much key/query budget is necessary and sufficient to represent its edges across arbitrary contexts.
In‑Context Learning and Algorithmic Views of Self‑Attention
A complementary line of theory frames Transformers—and attention in particular—as executing algorithms over the context. [38] analyze generalization and implicit model selection in in‑context learning; [65] give evidence that Transformers can implement gradient-descent‑like updates in context; and [20] characterize which simple function classes are learnable in context. These works clarify how attention can implement algorithmic behaviors over token sets, while our RGR focus quantifies the key–query capacity required to retrieve relational edges reliably.
Connectivity Patterns vs. Capacity in the Key/Query Channel
An alternative way to constrain attention is by controlling the connectivity pattern of the attention graph. Even $O(\ell)$-sparse patterns can be universal under appropriate designs [71], and systematic pruning of dense patterns maps out cost–performance frontiers [67]. Our analysis treats connectivity as not the bottleneck: given the ability to attend broadly, the limiting factor for RGR is how much relational information can be encoded and separated in keys/queries as $m$ and $m^{\prime}$ grow.
Head Specialization, Pruning, and Information Bottlenecks
Mechanistic interpretability consistently finds that specific heads specialize to linguistic relations [9, 63]. At the same time, many trained heads can be pruned with small accuracy loss [45, 64], indicating redundancy. Information-bottleneck analyses at the head/layer level quantify such redundancy and attribution in both language and vision models [49, 27], and architectural proposals target representation bottlenecks [21]. Our results supply a capacity-theoretic backbone for these observations: for RGR, performance transitions are governed primarily by $D_{K}$; distributing $D_{K}$ across heads reduces interference between superposed relations, but overly small $d_{k}$ per head incurs rank limits—predicting both specialization and safe pruning regimes.
Attention as Associative Memory vs. Relational Addressing
Modern Hopfield networks are equivalent, in a precise sense, to attention updates and can store exponentially many patterns in the associative dimension with single-step retrieval [50]. FFN layers in Transformers have also been interpreted as key–value memories [22]. Our results complement this memory-centric view by isolating the addressing budget: how much key/query capacity is required to select the correct neighbors (edges) for arbitrary contexts. Together these views separate storage capacity from the cost of accurate retrieval/selection in the key–query channel.
Graph Transformers and Structural Encodings
Expressivity of graph Transformers is shaped by structural encodings and higher-order tokenization. SEG-WL analyses show that structural features (e.g., SPIS encodings) set the attainable expressivity ceiling and can be matched by simple Transformer variants [75]. Higher-order graph Transformers reach (or fall short of) $t$-WL power depending on whether explicit tuple indices and structural signals are provided [74]. In the vision setting, [29] prove that ViTs can learn spatial structure under appropriate conditions, resonating with our assumptions that near-orthogonal embeddings and structural signals determine how efficiently edges can be packed and recovered; given such signals, our $D_{K}$-based bounds become tight predictors of success.
Summary.
Across expressivity, connectivity, memorization, superposition, interpretability, memory equivalence, and graph structure, prior work identifies the ingredients that make attention powerful and the constraints that limit it. We contribute a capacity-centric bridge: a concrete relational task (RGR) in which the total key dimension $D_{K}$ is the critical budget, with lower and upper bounds tight up to logarithmic factors, a principled multi-head advantage, and empirical thresholds that align with constructive algorithms.